### v1

# Balancing Intensity and Focality in Directional DBS Under Uncertainty: A Simulation Study of Electrode Optimization via a Metaheuristic L1L1 Approach

Abstract
Background and Objective: As Deep Brain Stimulation (DBS) technology advances toward directional leads and optimization-based current steering, this study aims to improve the selection of electrode contact configurations using the recently developed L1-norm regularized L1-norm fitting (L1L1) method. The focus is in particular on L1L1’s capability to incorporate a priori lead field uncertainty, offering a potential advantage over conventional approaches that do not account for such variability.
Methods: Our optimization framework incorporates uncertainty by constraining the solution space based on lead field attenuation. This reflects physiological expectations about the volume of tissue activated (VTA) and serves to avoid overfitting. By applying this method to 8- and 40-contact electrode configurations, we optimize current distributions within a discretized finite element (FE) model, focusing on the lead field’s characteristics. The model accounts for uncertainty through these explicit constraints, enhancing the feasibility, focality, and robustness of the resulting solutions.
Results: The L1L1 method was validated through a series of numerical experiments using both noiseless and noisy lead fields, where the noise level was selected to reflect attenuation within VTA. It successfully fits and regularizes the current distribution across target structures, with hyperparameter optimization extracting either bipolar or multipolar electrode configurations. These configurations aim at maximizing focused current density or prioritize a high-gain field ratio in a discretized FE model. Compared to traditional methods, the L1L1 approach showed competitive performance in concentrating stimulation within the target region while minimizing unintended current spread, particularly under noisy conditions.
Conclusions: The L1L1 method provides a valuable tool for assisting specialists in optimizing electrode configurations for DBS. By incorporating uncertainty directly into the optimization process, we obtain a noise-robust framework for current steering, allowing for variations in lead field models and simulation parameters.
keywords:
Deep Brain Stimulation (DBS) , Directional Lead , Anterior Nucleus of Thalamus , Convex Optimization , Metaheuristics
[…]
[inst1]Computing Sciences Unit, Faculty of Information Technology and Communication Sciences, Tampere University, Tampere, Finland
1 Introduction
Deep Brain Stimulation (DBS) [1, 2] is a neuromodulatory therapy for managing movement disorders such as Parkinson’s disease [3, 4, 5, 6], essential tremor [7], refractory epilepsy [8, 9, 10], and dystonia [11], particularly in patients with insufficient response to pharmacological treatments [12]. The therapy delivers electrical stimulation to specific subcortical brain regions through implanted electrodes. Among the various parameters that influence clinical outcomes, the configuration of the stimulating electrodes plays a critical role, yet remains insufficiently standardized and not fully understood.
The central idea in determining an electrode configuration relies on predicting and controlling the regions of axonal activation induced by the lead [13, 14]. The extent of this activation depend on several factors, including tissue conductivity [15, 16], electrode geometry [17, 18], and the anatomical features of the surrounding brain structures [19]. Electrode configurations are typically classified by the number and polarity of active contacts within the stimulation target. In a monopolar (or unipolar) configuration, one or more cathodic contacts reside within the target structure, while the return path is an anode located at the implantable pulse generator, which is electrically distant and treated as a reference at infinity. This results in a broadly spherical electric field centered around the active contact(s). In contrast, bipolar configurations employ both cathodic and anodic contacts located within the brain, allowing for more localized current steering between adjacent contacts. Multipolar configurations generalize this further by utilizing multiple active contacts of both polarities within the lead, enabling finer control of the field shape and distribution.
Through deliberate selection of electrode configurations, clinicians can control which anatomical regions around the directional DBS (dDBS) leads are stimulated [20, 21, 18, 22, 23, 24, 25, 26]. This capability enables personalized shaping of the electric field in terms of spatial extent, intensity, and directionality, optimizing therapeutic outcomes while reducing off-target effects. However, the increasing complexity of modern lead designs has introduced a combinatorial explosion in possible configurations. This has driven significant research efforts into advanced electrode geometries and computational optimization strategies for electric field shaping [27, 28, 29, 30, 21, 31].
In this study, we focus on computational modeling of brain anatomy and optimization techniques to identify effective bipolar and multipolar electrode configurations for dDBS leads. We employ the L1-norm regularized L1-norm fitting (L1L1) method [32], a sparse optimization framework designed to yield focal stimulation profiles by simultaneously penalizing the control vector and fitting residuals using L1-norms. This formulation promotes sparsity in the solution space, enabling interpretable and spatially selective activation patterns. Moreover, the L1L1 method naturally incorporates uncertainty in the lead field model, increasing robustness to measurement noise and reducing the risk of overfitting. These properties make it well-suited for individualized DBS modeling, where anatomical variability and signal noise present significant computational and neurophysiological challenges.
To constrain the optimization problem, we define the fitting region based on the spatial attenuation of the lead field, which approximates the clinically relevant volume of tissue activated (VTA). This region corresponds to the subset of neural tissue where the electric field amplitude exceeds a threshold sufficient for neuronal activation. Regions outside this boundary are excluded from the fitting procedure, as stimulation effects there are negligible or uncertain. Employing the VTA as a constraint provides a computationally tractable surrogate for detailed multicompartment neuron models that simulate extracellular potentials and neuronal activation dynamics, which are often prohibitively expensive in high-dimensional optimization tasks [33, 34, 35].
To quantitatively assess the robustness of the L1L1 optimization framework under realistic conditions, we expand the study to introduce additive Gaussian noise to the lead field matrix, simulating possible modeling errors and measurement variability. The noise magnitude is calibrated to reflect plausible physiological and instrumentation uncertainties such as strong deformation/distortions. We then evaluate the L1L1 method against two established baseline methods: the Reciprocity Principle (RP) [36], which leverages reciprocity theory to infer optimal stimulation patterns, and Tikhonov Regularized Least Squares (TLS) [37, 38], which balances data fidelity and solution smoothness through L2-norm regularization. This comparative analysis elucidates the trade-offs in spatial precision, sparsity, and noise resilience among the methods, highlighting the advantages of the L1L1 approach for individualized electrode configuration optimization.
2 Materials and Methods
2.1 Head Model Discretization
This study employs finite elements (FE) to solve the linear forward problem derived from the complete electrode model [39] and to illustrate the interpolated distributions on the FE mesh [40, 41]. The FE mesh is constructed using an openly accessible T1-weighted MRI-based dataset [42], which includes MRI data from a healthy 48-year-old right-handed individual who serves as our representative subject. The brain volume is segmented using the FreeSurfer Software Suite [43], and subcortical structures are extracted from FreeSurfer’s Aseg atlas (Fig. 1(a)). The head model’s parcellation is based on a multicompartment conductivity distribution, comprising 18 cortical and subcortical segments. The element-wise constant electrical conductivity values for these regions were set according to [44], and the subcortical nuclei values were taken from [45, 46]. Using this curated data, an unstructured tetrahedral boundary-fitted FE mesh is generated using the Zeffiro Interface (ZI) [47], with a resolution of 3.0 mm (millimeters). The volume of the thalamus compartment was further refined through three nested refinements [48] with each tetrahedron subdivided into four congruent parts, resulting in tetrahedra with diameters ranging from 0.1725 mm to 0.375 mm. This process led to a mesh consisting of 4.4 million nodes and 25.4 million elements (Fig. 1(b)). We define the anterior nucleus of the thalamus (ANT) as the stimulation target, similar to ANT-DBS, a clinically established approach for treating epilepsy [8, 9].
(a) Segmented head model domain depicting main subcortical structures.
(b) Refined and boundary-fitted thalamus sub-compartment.
(c) Superior, coronal, and sagittal view of target domain. The interthalamic adhesion can be easily noticed.
(d) 8- and 40-electrode contacts, respectively.
Figure 1: (1(a)): A head model including skull, forebrain, midbrain, hindbrain structures as well as directional deep brain stimulation (dDBS) lead positioned at the right-hemispheric anterior nucleus of thalamus (ANT). (1(b)): Refined and boundary-fitted thalamus sub-compartment with the distribution of tetrahedral mesh edges and an interpolated current intensity distribution shown on the surface. (1(c)): Placement of a 40-contact dDBS lead with a bipolar electrode configuration (anodal as red, and cathodal as blue). The dipolar target current (red) is depicted by the red pin, indicating the target position and orientation; a 6.0 millimeter (mm) region of interest (ROI) is displayed as an arrangement of source points (yellow squares) for High-Resolution Lead Field (HR-LF) matrix. (1(d)): Probing strategy of the 8- and 40-electrode contacts inspired by Abbott’s St. Jude Medical Infinity™ Directional Lead 6180 model and Medtronic-Sapiens, respectively.
2.2 Lead Field Matrix
The linear forward mapping can be formulated as a real $N\times K$ lead field matrix $\hat{\bf L}$, which relates current injections at the electrodes to the resulting volumetric current density distribution across different electrode configurations. This matrix characterizes the numerical relationship between the real $K\times 1$ current pattern ${\bf y=(y_{1},\ldots,y_{K})}$ and the real $N\times 1$ discretized volume current density $\hat{\bf x}=(\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{N})$. The latter variable represents an arrangement of uniformly distributed source points within the target domain, each with three degrees of freedom (DOFs) corresponding to three dipolar unit currents with Cartesian orientations.
We generate a lead field matrix for two different resolution levels: (i) Low-Resolution Lead Field (LR-LF), which corresponds to a target set of 244 $\times$ 3 spatial DOFs that cover the entire thalamus compartment. This robust interpretation offers a functional assessment of the model, serving as a benchmark for our numerical experiments. To validate whether our minimal interpretation is adequate, we compare the volumetric current density distribution with a locally detailed lead field matrix, (ii) High-Resolution Lead Field (HR-LF). This matrix consists of 3620 $\times$ 3 DOFs distributed across the thalamus, meaning that the mutual distance between nearest neighboring DOFs is reduced by 40% compared to LR-LF. Due to the increased computational cost of using a larger lead field size, our examination of HR-LF is confined to a spherical region of interest with a 6.0 mm radius, co-centered with the lead model (Fig. 1(c)).
2.3 Probing Strategy
[…]
The maximum total current injection supplied by all the active electrodes is limited to 4.0 mA (milliamperes) (5), and the maximum current per electrode contact is set to $\pm$ 2.0 mA (6). This can be considered a moderately typical setting, for example, 3.0 mA as a reference for monopolar patterns in [50]. The total absolute sum of electric potential corresponding to the contacts in the lead is normalized to meet the intended maximum current injection value (7). As a reference threshold for neural activation, we consider the current amplitude of 3.85 A/m2 (amperes per square meter) [51], which falls within the excitation current thresholds of 2.5 to 6.0 A/m2 estimated for nerve fibers in the upper limb area of the motor cortex at frequencies of 50.0 Hz and 2.44 kHz [52]. Our optimization framework maintains this reference threshold constant across all tested configurations. By holding temporal characteristics constant, we ensure that the observed variations in activation patterns arise from differences in spatial registrations and current steering, rather than from stimulation timing. This choice follows the framework established in previous studies by Deli et al. [53] and Schlaepfer et al. [54], where monopolar and bipolar stimulation configurations were evaluated using a consistent stimulation protocol, specifically maintaining a pulse width of 60 µs and a frequency of 130 Hz across all conditions. This design allowed the authors to isolate the effects attributable solely to electrode configuration by holding other key stimulation parameters constant, with amplitude as the only adjusted variable.
2.4 Optimization Algorithms
The aim of this study is the application of mathematical optimization algorithms to retrieve the current pattern ${\bf y}$ (4) within the local target region, aiming to optimally match the volumetric current distribution generated by the target current dipole modeling [55, 56, 51]. To quantify the optimal electrode configurations derived from the algorithm, we compare them based on the decision variables $\Gamma$, $\Xi$, and $\Theta$, as delineated by the optimization paradigm. The focused current density $\Gamma$ (A/m2) is defined as
${\Gamma}=\frac{{\bf x}_{1}^{T}{\bf L}_{1}{\bf y}}{\|{\bf x}_{1}\|_{2}}.$
The focality of the stimulation, or field ratio, $\Theta$, is defined as the ratio of the focused current density divided by the undesired nuisance current density ($\Xi$) and is defined as
${\Theta}=\frac{{\Gamma}}{\Xi}\quad\hbox{with}\quad\Xi=\frac{\|{\bf L}_{2}{\bf y}\|_{2}}{\sqrt{M}}.$
2.4.1 Bipolar Configurations
Bipolar configurations are derived through the reciprocity principle (RP) [36]. The principle states that the maximum focused current density, i.e., $\Gamma_{\hbox{\scriptsize{max}}}$, of an applied linear system can be achieved by limiting the electrode configuration to a pair of anodal and cathodal electrode contacts (Fig. 3). This configuration is obtained by multiplying the dipolar target current distribution by the transpose of the lead field matrix, selecting both the greatest positive and negative poles, and setting the remaining electrodes in the lead to zero. The reciprocity principle aims at maximizing current density at the expense of field ratio fidelity.
2.4.2 Multipolar Configurations
Multipolar configurations, i.e., leads with multiple active electrode contacts, are obtained via convex optimization techniques such as L1-norm regularized L1-norm fitting (L1L1) [32] and through Tikhonov-Regularized Least Squares (TLS) [37, 38] methods. Both methods employ a metaheuristic search strategy to explore only a carefully chosen fraction of an otherwise huge search space—trading guaranteed optimality for scalability and adaptability—to handle large, ill-behaved, or poorly understood problems within practical time and memory constraints [57]. The candidate solution is the one that can achieve the highest ratio between the (target) current density at the stimulation site and that in the surrounding adjacent areas (referred here to as nuisance current). The determination of this solution is defined by tuning a set of predefined hyperparameters along with the current pattern ${\bf y}$, subject to specific metacriteria, namely, maximizing the field ratio while ensuring $\Gamma\geq\Gamma_{0}>0$, thereby maintaining sufficient intensity at the target location (with $\Gamma_{0}=0.80$ mA in this case). Without a lower bound for intensity $\Gamma$, the intensity of the $\Theta$-maximizer is likely to diminish. A rough upper bound is provided as a L1-norm fitting threshold to suppress nuisance current density $\Xi$. In contrast, traditional least-squares formulations tend to produce smoothly distributed currents, which may be clinically suboptimal, as focal and sparse activation is often preferred to minimize side effects and enhance therapeutic precision. Moreover, least-squares approaches are sensitive to noise and outliers, often yielding non-robust solutions [58]. Detailed formulation of both the L1L1 and TLS methods is provided in A.
2.5 Uncertainty Modeling
Uncertainty refers to various physiological and technical factors that directly or indirectly affect the accuracy and reliability of the stimulation, including imprecision in lead placement, patient-specific anatomical variation, and inhomogeneities in tissue conductivity. In this study, all sources of uncertainty affecting the optimization problem are modeled as additive perturbations to the lead field. While this Gaussian noise model is a simplification, it enables a tractable first-step analysis of robustness. We acknowledge that in practice, the true uncertainty arises from a variety of unrelated sources.
[…]
(a) Left: The set $\mathcal{V}_{\delta}$ corresponding to PSNR $\delta^{-1}$ following from the uncertainty factors is shown by the blue dashed circle with radius $d_{\delta}$. The target is at the distance $d_{T}$ from the probe. Between the target and the boundary of $\mathcal{V}_{\delta}$ is a range $d_{\varepsilon}$ in which the lead field decays by the factor $\varepsilon$, which approximates the maximum achievable dynamic range $\varepsilon$ between the focused current density at the target position and in its surroundings. Right: One-dimensional interpretation of the lead field decay.
$\mathcal{V}_{-10\text{ dB}}$
$\mathcal{V}_{-20\text{ dB}}$
$\mathcal{V}_{-30\text{ dB}}$
$\mathcal{V}_{-40\text{ dB}}$
(b) The set $\mathcal{V}_{\delta}$ illustrated for the volumetric model of this study with respect to PSNR $\delta^{-1}$ levels from 10 to 40 dB.
Figure 2: (2(a)): A schematic of how the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ relates to the set $\mathcal{V}_{\delta}$ and to the approximate maximum dynamic range $\varepsilon^{-1}$ of the volumetric stimulation current field. (2(b)): The set $\mathcal{V}_{\delta}$ shown for PSNR 10–40 dB, providing a rough approximation of the maximum feasible coverage of volume of tissue activated (VTA) for an activation threshold set by $\delta$.
8-contact Electrode Configuration
Refer to caption RPRefer to captionRefer to caption Refer to caption L1L1(A)Refer to captionRefer to caption Refer to caption L1L1(B)Refer to captionRefer to caption Refer to caption TLSRefer to captionRefer to caption
40-contact Electrode Configuration
Refer to caption RPRefer to captionRefer to caption Refer to caption L1L1(A)Refer to captionRefer to caption Refer to caption L1L1(B)Refer to captionRefer to caption Refer to caption TLSRefer to captionRefer to caption
Figure 3: Examples of current injection patterns obtained using the Reciprocity Principle (RP), metaheuristic L1-norm regularized L1-norm fitting (L1L1) method under two different nuisance current thresholds (L1L1(A) with $\varepsilon=\{0:-160\}$ and L1L1(B) with $\varepsilon=\{0:-10\}$), and the Tikhonov-Regularized Least Square (TLS) method. The target position and orientation (parallel to the lead) is indicated by the red upward pin. The bar plots display the resulting electrode current amplitudes, while the contact maps visualize their spatial distribution over the electrodes; cathodal (blue), anodal (red), and inactive (gray/green). The RP method yields a bipolar configuration maximizing focused current density $\Gamma$, while the remaining methods searches for a multipolar configuration maximizing focality $\Theta=\Gamma/\Xi$ under an a priori constraint $\varepsilon$ on the nuisance current density $\Xi$, determined by the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ and the lead field attenuation. The L1L1(B) result reflects more strongly constrained $\Xi$ (smaller dynamic range), activating fewer contacts whereas L1L1(A) and TLS allow higher $\Theta$ values and broader excitation patterns. The voltage/current on the electrode contacts are ±$\pm 2.0$ mA (milliamperes).
We postulate that the volumetric current density distribution can be optimized for those degrees of freedom indexed by $i$, where the L2-norm (2-norm) $\|{\bf\tilde{l}}_{i}\|_{2}$ of the corresponding reduced lead field matrix row remains above the threshold. This row, denoted as ${\bf l}_{i}$, is first reduced (as described in Appendix 9) and then zero-averaged, taking the form
$\tilde{\bf l}_{i}=\left(L_{i,1}-\overline{L}_{i},L_{i,2}-\overline{L}_{i},\dots,L_{i,K}-\overline{L}_{i}\right),$
where $\overline{L}_{i}=K^{-1}\sum_{j=1}^{K}L_{i,j}$ represents the mean of the row entries. The threshold for optimization is determined by a peak signal-to-noise ratio (PSNR) constraint, ensuring that only degrees of freedom with a sufficiently strong contribution to the lead field are considered in the optimization process. We define the feasible set of degrees of freedom as
$\mathcal{V}_{\delta}=\{i\,|\,\|{\bf\tilde{l}}_{i}\|_{2}\geq\delta\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}\},$
where $\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}$ is the maximum L2-norm of the zero-averaged lead field rows. The parameter $\delta$ sets a relative threshold, filtering out degrees of freedom whose contributions are too weak to be considered meaningful. If $\delta$ is chosen to represent neural activation threshold, then $\mathcal{V}_{\delta}$ serves as an approximation of the maximum VTA (Fig. 2(a)).
Since lead field attenuation and uncertainty impose fundamental limits on stimulation accuracy, the PSNR constraint defines the theoretical limit of the optimization performance. If the target stimulation site corresponds to the $\ell$-th index, then the maximum achievable dynamic range $\varepsilon^{-1}$ between the current density at the target site and its surroundings is approximately
$\varepsilon\geq\delta\,\|{\bf\tilde{l}}_{\ell}\|_{2}^{-1}\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}.$
This relationship indicates that the minimum contrast in current density between the target location and surrounding regions is constrained by $\delta$ and the structure of the lead field matrix (Fig. 2(b)).
2.6 Numerical Experiments
Our numerical experiments are designed to evaluate the performance of the proposed optimization methods with respect to: (i) simulated 8-contact and 40-contact DBS lead configurations, (ii) noiseless and noisy forward mappings, (iii) orientation of the current dipole in the model (parallel or perpendicular with respect to the lead array) [17], and (iv) algorithm efficiency. The hardware used for these experiments includes a Dell Precision 5820 Workstation with 256 GB of RAM, a 10-core Intel i9–10900X CPU, and an NVidia Quadro RTX 4000 GPU. The GPU (graphics processing unit) was utilized to accelerate both forward and inverse computations.
Table 1: Summary of the proposed optimization methods. While the RP provides a bipolar reciprocal of the target, both L1L1 and TLS constitute a metaheuristic search over a two-dimensional hyperparameter lattice. L1L1 is subdivided into L1L1(A) and L1L1(B) based on two alternative a priori assumptions of the feasible $\varepsilon$ threshold range which is [-160, 0] and [-10, 0] dB, respectively. The regularization and the L2-norm weighting parameter intervals are chosen as in [32]. CPU time refers to the computing time needed to evaluate a single optimizer with fixed hyperparameter values for HR-LF with the 40-contact lead configuration.
Hyperparameter CPU
Tuning Range Role Time
Method Search Type (dB) (s)
RP Reciprocal 5.7E-03
L1L1(A) Metaheuristic $\alpha\in[-100,-30]$ Regularization 6.0
$\varepsilon\in[-160,0]$ Threshold
L1L1(B) Metaheuristic $\alpha\in[-100,-30]$ Regularization 3.7
$\varepsilon\in[-10,0]$ Threshold
TLS Metaheuristic $\gamma\in[-200,-110]$ Regularization 5.0E-02
$\beta\in[-50,40]$ Weight
The L1L1 method is examined using two alternative a priori assumptions for the feasible $\varepsilon$ threshold range: [-160, 0] dB and [-10, 0] dB, respectively. We refer to these settings as ”L1L1(A)” and ”L1L1(B)”. The former setting neglects the effect of uncertainty on the field ratio $\Theta$, while the second assumes that uncertainty is included within 10 dB dynamic range. As for TLS method, the regularization parameter $\gamma$ and the L2-norm weighting parameter $\beta$ intervals are chosen as in [32] (Table 1).
Gaussian noise is added to the lead field to examine spatial sensitivity; the effect of noise is assessed in an overall setting, where the candidate solution is determined for each target DOF of LR-LF under PSNR 40.0 dB noise corruption. While anatomical variations may cause spatially correlated errors, the implementation of noise offers a generic and tractable way to assess noise-robustness when the exact structure of uncertainty is unknown. It serves as a standard baseline for evaluating sensitivity to model perturbations.
The noise sensitivity is examined for two specific target positions (I) and (II) (Fig. 4) using PSNRs of 40.0 and 50.0 dB, along the HR-LF. When interpreted in terms of $\mathcal{V}_{\delta}$ (Figure 2(b)), PSNR 40.0 dB (50.0 dB) corresponds to the assumption that each possible target position in the set $\mathcal{V}_{-30\text{ dB}}$ ($\mathcal{V}_{-40\text{ dB}}$) can be stimulated with at least a 10 dB dynamic range $\varepsilon^{-1}$ for $\Theta$. This assumption aligns well with the $\varepsilon$ range of $[-10,0]$ selected for L1L1(B).
Figure 4: Visualization of the two synthetic stimulation targets, Target (I) and Target (II), analyzed separately in numerical experiments. Both targets are positioned within the right thalamus near a directional DBS lead. The figure also displays the lead’s center point and a concentric spherical region of interest (ROI) with 6 mm radius, representing an approximate VTA. This ROI defines the region where maximizing stimulation intensity while minimizing current spread to surrounding areas is considered feasible. As targets (I) and (II) are not right next to the probe, lead field uncertainty affects the optimization process. Bilateral thalamic structures are shown in gray for anatomical reference.
LR-LF HR-LF
Focused $\Gamma$ (A/m2) Nuisance $\Xi$ (A/m2) Ratio $\Theta$ (rel.) Focused $\Gamma$ (A/m2) Nuisance $\Xi$ (A/m2) Ratio $\Theta$ (rel.)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 5: Volumetric optimization results for the 8-contact probe and noiseless lead fields. Decision variable $\Gamma$, $\Xi$ and $\Theta$ for optimizers found via RP, TLS, L1L1(A), and L1L1(B) method illustrated as a function of the target position and orientation (parallel or perpendicular w.r.t. the probe). Left: distributions for a (i) Low-Resolution Lead Field (LR-LF) where the spatial degrees of freedom (DOFs) cover both thalamus lobes. Right: distribution for a (ii) High-Resolution Lead Field (HR-LF) where the DOFs is confined to a region of interest within a radius of 6.0 mm (millimeters) from the center of the lead. The optimal electrode configuration found by the L1L1 method maximizes the field ratio $\Theta$ subject to $\Gamma>=\Gamma_{0}$. The $\Theta$ distribution found using L1L1(A) is elevated as compared to the case of RP and TLS. The limited dynamic range ($\varepsilon$ range) of L1L1(B) is reflected in an overall slightly lower $\Theta$.
LR-LF HR-LF
Focused $\Gamma$ (A/m2) Nuisance $\Xi$ (A/m2) Ratio $\Theta$ (rel.) Focused $\Gamma$ (A/m2) Nuisance $\Xi$ (A/m2) Ratio $\Theta$ (rel.)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 6: Volumetric optimization results for the 40-contact probe and noiseless lead fields. As compared to the results obtained with 8-contact probe the overall level of $\Theta$ is greater for L1L1(A), TLS and RP, while for L1L1(B) with limited dynamic range ($\varepsilon$ range) there is no clear increase.
3 Results
The applied algorithms generate distinct distributions of the variables $\Gamma$, $\Xi$, and $\Theta$, which can be differentiated based on how they are spatially organized across the target domain under the optimal electrode configurations for the 8- (Fig. 5) and 40-contact (Fig. 6) leads. The direction of the dipole current plays a key role in shaping electric field targeting. The anode–cathode pair exhibiting the strongest positive and negative values may activate electrodes located as close as $0.189\pm 0.009\,\text{mm}$ or as far as $6.375\pm 0.375\,\text{mm}$ apart across the lead. This effect becomes evident when comparing the corresponding distributions: in the parallel orientation, the targeting is highly concentrated near the tip of the probe, whereas the perpendicular orientation results in a broader spatial spread, with activity shifted medially relative to the lead.
The bipolar configurations obtained via RP yield the overall highest focused current density $\Gamma$; however, this comes at the expense of focality, as indicated by elevated nuisance current $\Xi$ and a reduced field ratio $\Theta$. This trade-off is understandable, as the multipolar configurations are designed to maximize the field ratio $\Theta$ subject to the constraint $\Gamma\geq\Gamma_{0}>0$, with $\Gamma_{0}=0.8\,\text{mA}$, thereby allowing activation of more than two contacts. In the noiseless setting, the L1L1(B) reconstructions with $\varepsilon\in[-10,0]\,\text{dB}$ yield slightly higher values of $\Gamma$ and $\Xi$, but exhibit a more moderate $\Theta$ compared to L1L1(A) with $\varepsilon\in[-160,0]\,\text{dB}$. Nonetheless, both L1L1(A) and L1L1(B) demonstrate superior nuisance current suppression relative to TLS and RP.
Figure 7 and Table 2 compare the noiseless and noisy ($\text{PSNR}=40\,\text{dB}$) results for $\Gamma$ and $\Theta$ obtained using L1L1(A) and L1L1(B) and LR-LF. The results suggest that, unlike L1L1(A), L1L1(B) is more robust to lead field noise, yielding similar outcomes in both settings. In contrast, L1L1(A), which uses a smaller lower bound for $\varepsilon$, tends to overfit the noiseless data. This leads to unrealistically high values of the field ratio $\Theta$: up to $27.07$ in the 8-contact configuration and as high as $536.05$ in the 40-contact case. However, in the presence of noise, the results produced by L1L1(A) fall within a similar range to those of L1L1(B), indicating that the overfitting is specific to the noiseless scenario.
Figure 8 further examines the noise sensitivity of RP, TLS, L1L1(A), and L1L1(B) for two specific target positions, illustrated in Figure 4, covering a sample of twenty different noise realizations for both PSNR 40 and 50 dB. The results demonstrate that in a noisy situation, L1L1(A) and L1L1(B) provide similar distributions for $\Gamma$ and $\Theta$, hence, also for $\Xi$. Compared to RP and TLS, the L1L1 approach seems to provide a somewhat elevated field ratio $\Theta$ with a difference that is only marginal or absent with PSNR 40 dB, but observable with PSNR 50 dB. Accordingly, the potential benefit of using L1L1 depends critically on the level of modeling uncertainty, and it vanishes when this uncertainty becomes excessive. Of the four optimization approaches tested, L1L1(B) provides the closest agreement between noiseless and noisy outcomes and proves particularly effective in maximizing the field ratio $\Theta$ in the presence of noise.
Table 2: The maximum values of the decision variable (DV) $\Theta$ (field ratio) found by L1L1(A) and L1L1(B) searches and the corresponding values of $\Gamma$ (focused current density) for the low-resolution lead field (LR-LF) and its target set (Figure 7) covering both parallel and perpendicular target orientations of the DBS lead in both noiseless and noisy (PSNR 40 dB) setting.
8-contacts Multipolar
L1L1(A), $\varepsilon\in[-160,0]$ L1L1(B), $\varepsilon\in[-10,0]$
DV Orientation No Noise PSNR 40 dB No Noise PSNR 40 dB
$\Gamma$ Parallel 2.42 2.27 2.27 2.27
Perpendicular 3.04 2.35 2.28 2.34
$\Theta$ Parallel 21.53 5.56 13.08 5.48
Perpendicular 27.07 7.49 13.18 7.22
40-contacts Multipolar
L1L1(A), $\varepsilon\in[-160,0]$ L1L1(B), $\varepsilon\in[-10,0]$
DV Orientation No Noise PSNR 40 dB No Noise PSNR 40 dB
$\Gamma$ Parallel 2.20 1.74 2.02 2.24
Perpendicular 2.24 2.29 3.37 2.23
$\Theta$ Parallel 526.44 7.56 9.17 8.49
Perpendicular 536.05 9.96 10.97 8.47
8-contact Electrode Configuration
  Noiseless lead field Refer to caption A1. Focus $\Gamma$ (A/m2)     Refer to caption B1. Ratio $\Theta$ (rel.) Noisy lead field Refer to caption C1. Focus $\Gamma$ (A/m2)     Refer to caption D1. Ratio $\Theta$ (rel.)
40-contact Electrode Configuration
  Noiseless lead field Refer to caption A2. Focus $\Gamma$ (A/m2)     Refer to caption B2. Ratio $\Theta$ (rel.) Noisy lead field Refer to caption C2. Focus $\Gamma$ (A/m2)     Refer to caption D2. Ratio $\Theta$ (rel.)
L1L1(A), $\varepsilon\in[-160,0]$, Parallel
L1L1(B), $\varepsilon\in[-10,0]$, Parallel
L1L1(A), $\varepsilon\in[-160,0]$, Perpendicular
L1L1(B), $\varepsilon\in[-10,0]$, Perpendicular
Figure 7: Performance of L1L1(A) and L1L1(B) for optimizing the field ratio $\Theta$ subject to criterion $\Gamma\geq\Gamma_{0}$ with noiseless and noisy (PSNR 40 dB) lead field. L1L1(A) and L1L1(B) suppress the nuisance current density $\Xi$ pattern using two different dynamic ranges, namely, $\varepsilon\in[-160,0]$ (blue) and $\varepsilon\in[-10,0]$ (red), respectively. The results are sub-grouped by the direction of the current dipole, whether its alignment is parallel (solid line) or perpendicular (dotted line) with the applied lead. The degrees of freedom have been sorted according to ascending field ratio $\Theta$. The results show that the overly optimistic dynamic range applied in L1L1(A) results in very high field values of $\Theta$ in the noiseless case, which are, however, absent in the noisy one. In contrast, L1L1(B) is only marginally affected by the noise due to more realistic assumption of the maximum obtainable dynamic range under uncertainty.
8-contact Electrode Configuration
  Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (I) Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (II)
40-contact Electrode Configuration
  Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (I) Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (II)
Figure 8: Boxplots showing the focused current density $\Gamma$ and field ratio $\Theta$ obtained with a noise-corrupted HR-LF and the 8-contact electrode configuration for test targets (I) and (II). The optimizers have been found through RP, TLS, L1L1(A), and L1L1(B) search. Each bar collates a sample of twenty decision variable values corresponding to optimizers obtained with as many independent lead field noise realizations (PSNR 40 or 50 dB). The box shows the interquartile range (IQR) between the first quantile $Q1$ and the third one $Q3$, with a horizontal line at the median. The whiskers extend from the edges of the box to the smallest and largest values within $[Q1-1.5\times\text{IQR},Q3+1.5\times\text{IQR}]$, where $\text{IQR}=Q3-Q1$. Data points outside this range are considered outliers and are plotted individually as markers. The results obtained with L1L1(A) and L1L1(B) demonstrate a comparable sensitivity under lead field noise, suggesting that the limited dynamic range of L1L1(B) does not downgrade the L1L1 method’s performance in situations involving uncertainty. Furthermore, both L1L1(A) and L1L1(B) find an overall elevated field ratio compared to RP and TLS with the greater PSNR level 50 dB, which allows for an improved +10 dB dynamic range for the optimization algorithm as compared to PSNR 40 dB.
4 Discussion
This study explores the application of the recently introduced metaheuristic L1L1 method [32], to optimize the current patterns for DBS leads. The core focus lies in L1L1’s capacity to incorporate a priori uncertainty through an explicit constraint, offering a notable advantage over traditional approaches, which often fail to account for such variability [37, 38, 36, 59]. Specifically, we define the uncertainty through a threshold on the nuisance current density, which is determined by the relative attenuation of the lead field with respect to a target site. This attenuation is quantified in decibels and corresponds to typical falloff patterns observed in VTA. By reducing the dependence on conventional parameter tuning methods, such as trial-and-error [60, 61], this approach paves the way for more sophisticated techniques to make more informed decisions, even in the presence of inherent uncertainties in DBS settings and simulations [62, 63, 64, 65].
Optimization was carried out using three parameters: the focused current density $\Gamma$, the nuisance current density $\Xi$, and field ratio $\Theta=\Gamma/\Xi$. To assess robustness and focality under uncertainty, the recently introduced L1L1 method was applied in two variants, L1L1(A) and L1L1(B), differing in the range of the nuisance current threshold $\varepsilon$, which reflects varying levels of a priori modeling uncertainty. Specifically, L1L1(B) constrains $\varepsilon$ to the interval $[-10,0]$ dB, accounting for higher uncertainty, whereas L1L1(A) utilizes a broader range $[-160,0]$ dB. For comparison, the Reciprocity Principle (RP) and Tikhonov-Regularized Least Squares (TLS) methods were included as reference approaches.
The results reveal that the spatial distributions of $\Gamma$, $\Xi$, and $\Theta$ are strongly influenced by both the optimization method and the orientation of the dipole current. Parallel configurations resulted in highly localized activation near the tip of the lead, whereas perpendicular configurations yielded a broader, medially shifted targeting. A comparative analysis across varying noise levels demonstrated that L1L1 methods consistently outperformed RP and TLS in maintaining higher and more stable values of $\Theta$, especially at higher PSNR values.
In noiseless conditions, solutions obtained with a low nuisance threshold $\varepsilon\in[-160,0]\,\text{dB}$, as in L1L1(A), often exhibited unrealistically high field ratio values $\Theta$, sometimes up to fifty times greater than those derived using the more conservative threshold $\varepsilon\in[-10,0]\,\text{dB}$ in L1L1(B). This tendency suggests overfitting when the model is not constrained by uncertainty. Introducing this noise regularized the solution space, preventing the algorithm from converging to over-idealized solutions that mirrored the original model structure. These findings underscore the significance of explicitly modeling uncertainty and justify the constraint imposed on the nuisance current in L1L1(B), which leads to more realistic and generalizable solutions. Overall, these observations highlight the potential of L1L1(B) as a robust and uncertainty-aware optimization strategy for dDBS lead configuration.
The estimates for the set $\mathcal{V}_{\delta}$, where the lead field amplitude attenuates by a factor $\delta$ for $\delta\in[-40,-10]$ dB, align closely with established knowledge regarding VTA. Specifically, for $\delta\approx-40\,\text{dB}$, our model predicts attenuation within a spatial region consistent with a VTA radius of approximately 4.0 mm, in line with the findings of Butson’s [33] and McIntyre’s [20]. This agreement suggests that the noise levels used, with PSNR values of 40 dB and 50 dB, represent reasonable bounds for physiological uncertainty. At 40 dB, the dynamic range of $\Theta$ is determined by the maximum lead field attenuation within the VTA, while 50 dB further supports the assumption that each position within the VTA can be independently and coherently targeted. In this context, the L1L1 method, particularly L1L1(B), demonstrated superior targeting capabilities compared to RP and TLS, especially under noisy conditions.
When comparing the electrode configurations, increasing the number of contact points generally enhances both adaptability and control over the resulting electric field. This improvement is driven by several factors, including reduced distances between electrodes and target locations within the lead field matrix, better alignment of the induced electric field with the target current direction, and a greater capacity for flexible fitting due to the increased number of degrees of freedom.
The findings of this study provide a proof-of-concept for integrating a priori uncertainty modeling into numerically simulated dDBS optimization. When relevant sources of uncertainty, such as anatomical variability, conductivity inhomogeneities, or lead placement deviations, are known or can be estimated, the L1L1 method might be employed to improve the focality and directionality of stimulation, enabling more selective targeting of desired structures while reducing activation in surrounding regions. In clinical practice, such improvements might lead to more personalized and effective treatment strategies, enhancing therapeutic precision, and minimizing side effects [10].
4.1 Limitations and Future Work
While the current uncertainty modeling strategy provides a principled framework for constraining the optimization based on lead field attenuation, it also has several limitations. First, the threshold values used to define acceptable nuisance current levels are heuristic and may not fully capture the complexity of individual patient anatomies or tissue conductivities. This simplification assumes spatial uniformity in attenuation, which may not hold, for example, in the presence of anisotropic or heterogeneous tissue properties [44, 66]. Additionally, the approach relies on a static surrogate model of the VTA and does not directly incorporate a probabilistic model of uncertainty [62], which would allow for a more comprehensive treatment of variability. Potential alternative surrogate approaches exist, such as, threshold-based curve fitting [67], artificial neural networks [68], and statistical emulators such as Gaussian process classifiers [69, 70], which approach VTA in different ways.
Compared to the reference methods of this study, the L1L1 method is also slightly more demanding in terms of computing time. However, the increase is not significant and does not limit its practical use.
Other mathematical optimization techniques, such as Tikhonov (TLS) [38, 37], may offer promising alternatives alongside the L1L1 method for further refining the a priori uncertainty modeling adopted in this study. A logical direction for future research is the development of a posteriori models, where the current approach could be validated in more advanced experimental or clinical settings. This would enable a deeper understanding of the potential hyperparameter realizations in real-world applications. Translating these computational findings into clinical practice will require rigorous validation through empirical testing and clinical trials.
An ongoing work is the inclusion of more patient-specific anatomical and physiological data, e.g., a thalamic segmentation, to ensure more personalized and accurate stimulation strategies. Therefore, we explore the use of FE-based discretizations incorporating thalamic subdivisions via boundary-fitted multicompartment FE meshing approach and a complete electrode model, as utilized in this study.
5 Conclusions
This study represents an advancement in optimizing the volumetric current fields for DBS by leveraging the metaheuristic L1L1 method. Previous studies on contact selection have used convex optimization and patient-specific models to determine optimal stimulation settings. These methods perform well when all input parameters are fixed. However, they do not account for uncertainties in factors such as patient anatomy, lead placement, and tissue properties, which often vary in clinical practice. In principle, our L1L1-based method can address these uncertainties. Its main advantage is the ability to target the desired area while avoiding overfitting to noisy data.
Our findings demonstrate the method’s ability to refine electrode contact point configurations under varying levels of a priori lead field uncertainty. By achieving relatively high field ratios with a carefully chosen $\varepsilon$ value, the L1L1 method shows promise in enhancing the precision and effectiveness of DBS stimulation within uncertainty constraints. This, in turn, has the potential to improve individualized stimulation outcomes. An ongoing work is to integrate patient-specific data, such as thalamic segmentation, to improve stimulation accuracy. This involves using boundary-fitted multicompartment FE meshes and a complete electrode model, as in this study.
6 Conflict of Interest
The authors certify that this study is a result of purely academic, open, and independent research. They have no affiliations with or involvement in any organization or entity with a financial interest or non-financial interest, such as personal or professional relationships, affiliations, knowledge, or beliefs in the subject matter or materials discussed in this manuscript.
7 Author Contributions
FGP: Conceptualization, Formal analysis, Investigation, Methodology, Software, Visualization, Writing—original draft. AL: Conceptualization, Data curation, Investigation, Methodology, Project administration, Writing—original draft. MS: Formal analysis, Supervision, Validation, Writing—review & editing. SP: Conceptualization, Formal analysis, Funding acquisition, Investigation, Methodology, Project administration, Resources, Software, Supervision, Validation, Visualization, Writing—review & editing.
8 Funding
FGP, MS, and SP were supported by the Research Council of Finland through the Center of Excellence in Inverse Modelling and Imaging 2018–2025 (353089); the researcher exchange (DAAD) project ”Non-invasively reconstructing and inhibiting activity in focal epilepsy” (354976, 367453); the ERA PerEpi project (344712), and Flagship of Advanced Mathematics for Sensing, Imaging and Modelling (FAME) (359185, 359198)
9 Acknowledgments
The authors would like to thank Prof. Dr. rer. nat. Carsten H. Wolters, researchers and clinicians from the Institute for Biomagnetism and Biosignalanalysis (IBB), and the PerEpi consortium for their continuous support, discussions, and feedback regarding non-invasive brain stimulation topics, regression analysis, and the seminars prepared during the elaboration of this study.
Appendix A Mathematical Optimization Scheme
We consider a weighted optimization scheme governed by the linear system
${\bf L}{\bf y}={\bf x},$ (1)
where $\mathbf{y}\in\mathbb{R}^{K}$ is the current pattern to be optimized, $\mathbf{x}\in\mathbb{R}^{N}$ is the discretized volume current density vector, and $\mathbf{L}\in\mathbb{R}^{N\times K}$ is the lead field matrix defining a linear forward mapping from current patterns to the resulting current density field. This system is derived by component-wise splitting of the full lead field matrix $\hat{\mathbf{L}}\in\mathbb{R}^{N\times K}$ and the full current field $\hat{\mathbf{x}}\in\mathbb{R}^{N}$, such that
$\hat{\mathbf{L}}=\begin{pmatrix}\hat{\mathbf{L}}_{1}\\ \hat{\mathbf{L}}_{2}\end{pmatrix},\quad\hat{\mathbf{x}}=\begin{pmatrix}\hat{\mathbf{x}}_{1}\\ \mathbf{0}\end{pmatrix},$ (2)
where $\hat{\mathbf{x}}_{1}\in\mathbb{R}^{N_{1}}$ corresponds to the target region of interest and the zero component enforces field suppression elsewhere. The reduced system (1) is then defined by
$\mathbf{L}=\begin{pmatrix}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{pmatrix}=\begin{pmatrix}\mathbf{P}\hat{\mathbf{L}}_{1}\\ \hat{\mathbf{L}}_{2}\end{pmatrix},\quad\mathbf{x}=\begin{pmatrix}\mathbf{x}_{1}\\ \mathbf{0}\end{pmatrix}=\begin{pmatrix}\mathbf{P}\hat{\mathbf{x}}_{1}\\ \mathbf{0}\end{pmatrix}.$ (3)
Here, $\mathbf{P}\in\mathbb{R}^{N_{1}\times N_{1}}$ is a projection matrix that projects vectors onto the direction of $\hat{\mathbf{x}}_{1}$, which is assumed to have fixed direction. This projection ensures that alignment between the induced and target fields is measured only along the direction of interest. In this formulation, the upper part $\hat{\mathbf{L}}_{1}\mathbf{y}=\hat{\mathbf{x}}_{1}$ defines the focused current field in the target region, while the lower part $\hat{\mathbf{L}}_{2}\mathbf{y}=\mathbf{0}$ defines the nuisance field to be suppressed outside the region of interest.
A.1 L1-norm regularized L1-norm fitting
The general form of the L1-norm regularized L1-norm fitting (L1L1) method described by Galaz Prieto et. al [32, 57] is an optimization problem designed to minimize a linear objective function of continuous real variables subject to linear constraints. The optimization problem needs to find the best matching between ${\bf y}$, and the focused field via ${\bf L}{\bf y}={\bf x}$. The task is
[…]
subject to $\quad{\bf y}\preceq\gamma{\bf 1},$ (5)
$\quad\|{\bf y}\|_{1}\leq\mu,$ (6)
$\quad\sum_{\ell=1}^{K}y_{\ell}=0,\,\quad\ell=1,\ldots,K.$ (7)
The linear constraint 5 implies that the individual current injection for each $\ell$-th electrode is limited to a $\Gamma$ value; constraint 6 implies the total current injection flowing through the system is within a safety $\mu$-value limit; and constraint 7 indicates the total sum of electrical activity from every active electrode in the system must be equal to zero. The regularization parameter ${\bf\alpha}$ sets the L1-regularization level regarding a scaling value $\zeta=\|{\bf L}\|_{1}$. The function
$\Psi_{\varepsilon}[{\bf w}]_{m}=\max\{\,|w_{m}|,\varepsilon\,\}\quad\hbox{for}\quad m=1,\ldots,M,$ (8)
where ${\bf w}=(w_{1},w_{2},\ldots,w_{M})$ sets the nuisance field threshold $0\leq\varepsilon\leq 1$ with respect to the scaling value $\nu=\|{\bf x}\|_{\infty}$, meaning that entries $({\bf L}_{2}{\bf y})_{m}$ with an absolute value below $\varepsilon\nu$ do not actively contribute to the minimization process due to the threshold. The constraint support index $\{\,m\,:\,|({\bf L}_{2}{\bf y})_{m}|\geq{\varepsilon\nu}\,\}$, i.e., sets the contributing to the value of the objective function.
A.2 Tikhonov-Regularized Least Squares
In TLS estimation [37, 38] is formulated as the following optimization problem:
$\min_{{\bf y}}\left\{\|{\bf L}_{1}{\bf y}-{\bf x}_{1}\|^{2}_{2}+\gamma^{2}\beta^{2}\|{\bf L}_{2}{\bf y}\|^{2}_{2}+\gamma^{2}\varsigma^{2}\|{\bf y}\|^{2}_{2}\right\},$ (9)
where $\varsigma=\|{\bf L}\|_{2}$. To enhance focality, the nuisance field weight $\beta\geq 0$ is treated as a variable parameter. The solution to (9) satisfies the linear system:
$\left({\bf L}_{1}^{T}{\bf L}_{1}+\gamma^{2}\beta^{2}{\bf L}_{2}^{T}{\bf L}_{2}+\gamma^{2}\varsigma^{2}{\bf I}\right){\bf y}={\bf L}_{1}^{T}{\bf x}_{1}.$ (10)
Both the regularization parameter and the nuisance field weight influence the above equations. However, different methods handle the latter distinctly: in L1L1, it appears as an explicit threshold, whereas in the TLS framework, it is incorporated as a weighted term, leading to a distinct dependence of the solution.
For a more detailed description of the methods above, forward model and comparative results using other optimization approaches such as the L1-norm regularized L2-norm fitting (L1L2) method, a variant of the L1L1 method, please consult the [32] paper. For technicalities and explanation of the mathematical optimization packages, algorithms and search methodologies applied, see the follow-up [57] paper.

### v2

@@
-Balancing Intensity and Focality in Directional DBS Under Uncertainty: A Simulation Study of Electrode Optimization via a Metaheuristic L1L1 Approach
+Balancing Intensity and Focality in Directional DBS Under Uncertainty: A Simulation Study of Electrode Optimization via Metaheuristic L1L1 Method
 Abstract
-Background and Objective: As Deep Brain Stimulation (DBS) technology advances toward directional leads and optimization-based current steering, this study aims to improve the selection of electrode contact configurations using the recently developed L1-norm regularized L1-norm fitting (L1L1) method.
+Background and Objective: As Deep Brain Stimulation (DBS) technology advances toward directional leads and optimization-based current steering, this study aims to improve the selection of electrode contact configurations using the recently developed metaheuristic L1-norm regularized L1-norm fitting (L1L1) method.
 The focus is in particular on L1L1’s capability to incorporate a priori lead field uncertainty, offering a potential advantage over conventional approaches that do not account for such variability.
@@
 keywords:
-Deep Brain Stimulation (DBS) , Directional Lead , Anterior Nucleus of Thalamus , Convex Optimization , Metaheuristics
+Deep Brain Stimulation (DBS) , Anterior Nucleus of Thalamus , Convex Optimization , Metaheuristics
 […]
-[inst1]Computing Sciences Unit, Faculty of Information Technology and Communication Sciences, Tampere University, Tampere, Finland
+[…]
 1 Introduction
-Deep Brain Stimulation (DBS) [1, 2] is a neuromodulatory therapy for managing movement disorders such as Parkinson’s disease [3, 4, 5, 6], essential tremor [7], refractory epilepsy [8, 9, 10], and dystonia [11], particularly in patients with insufficient response to pharmacological treatments [12].
-The therapy delivers electrical stimulation to specific subcortical brain regions through implanted electrodes.
-Among the various parameters that influence clinical outcomes, the configuration of the stimulating electrodes plays a critical role, yet remains insufficiently standardized and not fully understood.
-The central idea in determining an electrode configuration relies on predicting and controlling the regions of axonal activation induced by the lead [13, 14].
-The extent of this activation depend on several factors, including tissue conductivity [15, 16], electrode geometry [17, 18], and the anatomical features of the surrounding brain structures [19].
-Electrode configurations are typically classified by the number and polarity of active contacts within the stimulation target.
-In a monopolar (or unipolar) configuration, one or more cathodic contacts reside within the target structure, while the return path is an anode located at the implantable pulse generator, which is electrically distant and treated as a reference at infinity.
-This results in a broadly spherical electric field centered around the active contact(s).
-In contrast, bipolar configurations employ both cathodic and anodic contacts located within the brain, allowing for more localized current steering between adjacent contacts.
-Multipolar configurations generalize this further by utilizing multiple active contacts of both polarities within the lead, enabling finer control of the field shape and distribution.
-Through deliberate selection of electrode configurations, clinicians can control which anatomical regions around the directional DBS (dDBS) leads are stimulated [20, 21, 18, 22, 23, 24, 25, 26].
-This capability enables personalized shaping of the electric field in terms of spatial extent, intensity, and directionality, optimizing therapeutic outcomes while reducing off-target effects.
-However, the increasing complexity of modern lead designs has introduced a combinatorial explosion in possible configurations.
-This has driven significant research efforts into advanced electrode geometries and computational optimization strategies for electric field shaping [27, 28, 29, 30, 21, 31].
-In this study, we focus on computational modeling of brain anatomy and optimization techniques to identify effective bipolar and multipolar electrode configurations for dDBS leads.
-We employ the L1-norm regularized L1-norm fitting (L1L1) method [32], a sparse optimization framework designed to yield focal stimulation profiles by simultaneously penalizing the control vector and fitting residuals using L1-norms.
-This formulation promotes sparsity in the solution space, enabling interpretable and spatially selective activation patterns.
-Moreover, the L1L1 method naturally incorporates uncertainty in the lead field model, increasing robustness to measurement noise and reducing the risk of overfitting.
-These properties make it well-suited for individualized DBS modeling, where anatomical variability and signal noise present significant computational and neurophysiological challenges.
-To constrain the optimization problem, we define the fitting region based on the spatial attenuation of the lead field, which approximates the clinically relevant volume of tissue activated (VTA).
-This region corresponds to the subset of neural tissue where the electric field amplitude exceeds a threshold sufficient for neuronal activation.
-Regions outside this boundary are excluded from the fitting procedure, as stimulation effects there are negligible or uncertain.
-Employing the VTA as a constraint provides a computationally tractable surrogate for detailed multicompartment neuron models that simulate extracellular potentials and neuronal activation dynamics, which are often prohibitively expensive in high-dimensional optimization tasks [33, 34, 35].
-To quantitatively assess the robustness of the L1L1 optimization framework under realistic conditions, we expand the study to introduce additive Gaussian noise to the lead field matrix, simulating possible modeling errors and measurement variability.
-The noise magnitude is calibrated to reflect plausible physiological and instrumentation uncertainties such as strong deformation/distortions.
-We then evaluate the L1L1 method against two established baseline methods: the Reciprocity Principle (RP) [36], which leverages reciprocity theory to infer optimal stimulation patterns, and Tikhonov Regularized Least Squares (TLS) [37, 38], which balances data fidelity and solution smoothness through L2-norm regularization.
-This comparative analysis elucidates the trade-offs in spatial precision, sparsity, and noise resilience among the methods, highlighting the advantages of the L1L1 approach for individualized electrode configuration optimization.
+Deep Brain Stimulation (DBS) [volkmann2002introduction, Perlmutter_2006_DBS, Gardner_2013_history_DBS] is a neuromodulatory therapy used to alleviate symptoms in patients with movement and neurological disorders, including, e.g., Parkinson’s disease [rodriguez2005bilateral, picillo2016programming, krack2019deep, steffen2020bipolar], essential tremor [benabid1996chronic], refractory epilepsy [LEHTIMAKI_2016_DBS_targetting, Jarvenpaa_2020_Improving_ANTDBS, fasano2021experience], and dystonia [johnson2013neuromodulation], by implanting an electrode lead into a targeted subcortical region and delivering controlled electrical stimulation through its contacts to modulate pathological neural activity.
+The central objective when determining this region of axonal activation is to shape the resulting volume of tissue activated (VTA), elicited by activating specific electrode contacts along the lead, so that it aligns with the intended therapeutic target [mcintyre_2004_cellular, butson_2006_electrodesign, butson_2007_patient, miocinovic_2006_subthalamic].
+Computational representations of DBS activation patterns are often categorized by the number and polarity of active contacts.
+In monopolar configurations, one or more cathodic contacts are driven against the distant implantable pulse generator (IPG), which serves as a reference at infinity [volkmann_basic_algorithm_DBS_2006].
+Bipolar configurations confine both anode and cathode to contacts on the lead, creating a locally circulating current and producing a more spatially restricted field [steffen2020bipolar, DELI_2011_UnipolarBipolarComparison].
+Multipolar configurations further extend the programmable space by allowing multiple contacts with mixed polarities to shape the field in a more flexible manner.
+As contemporary leads incorporate segmented or high-density arrays, the resulting combinatorial growth in possible patterns renders exhaustive manual programming infeasible [Anderson_2018].
+Effective steering of the VTA depends on a combination of tissue properties [butson_2005_tissue, McIntyre_2006_OptimizingDBS], electrode geometry and orientation [willsie2015computational], and the anatomical variability of the target region [wei_2005_currentdensityDBS, van2015novel].
+Clinical investigations such as [contarino2014directional, pollo2014directional, dembek2017directional] and the synthesis in [Steigerwald_dDBS_2019] demonstrated directional steering empirically through stimulation mapping.
+Complementary computational studies, including [alonso2016investigation, zhang2020steering], examined how geometry, contact orientation, and heterogeneous tissue conductivities influence field distribution and VTA through finite element modeling.
+Although [Anderson_2018] illustrates how convex formulations can be incorporated into directional lead control, it represents one contribution within a broader set of emerging model-based strategies aimed at systematically optimizing stimulation patterns for directional DBS.
+We employ the metaheuristic L1-norm regularized L1-norm fitting (L1L1) method [galazprieto_2022_L1vsL2] to retrieve electrode configurations over standard 8-contact and a complex 40-contacts directional DBS leads.
+The L1L1 method treats both the control vector and the data misfit with L1 penalties, which makes the optimization resistant to noise and model mismatch while discouraging diffuse or overly complex solutions, unlike least-squares formulations.
+We include configurations following Reciprocity Principle (RP2) [FERNANDEZCORAZZA2020116403], which leverages reciprocity theory to infer optimal stimulation patterns, and Tikhonov Regularized Least Squares (TLS) [dmochowski2011optimized], which balances data fidelity and solution smoothness through L2-norm regularization.
+For clarification: the terms “L1” and “L2” refer to the discrete 1- and 2-norms, respectively, of finite-dimensional vectors.
+The capitalized form L1L1 is used solely as the algorithm name and does not refer to the Lebesgue functional spaces, i.e., $L^{1}$ or $L^{2}$.
 2 Materials and Methods
-2.1 Head Model Discretization
-This study employs finite elements (FE) to solve the linear forward problem derived from the complete electrode model [39] and to illustrate the interpolated distributions on the FE mesh [40, 41].
-The FE mesh is constructed using an openly accessible T1-weighted MRI-based dataset [42], which includes MRI data from a healthy 48-year-old right-handed individual who serves as our representative subject.
-The brain volume is segmented using the FreeSurfer Software Suite [43], and subcortical structures are extracted from FreeSurfer’s Aseg atlas (Fig.
-1(a)).
-The head model’s parcellation is based on a multicompartment conductivity distribution, comprising 18 cortical and subcortical segments.
-The element-wise constant electrical conductivity values for these regions were set according to [44], and the subcortical nuclei values were taken from [45, 46].
-Using this curated data, an unstructured tetrahedral boundary-fitted FE mesh is generated using the Zeffiro Interface (ZI) [47], with a resolution of 3.0 mm (millimeters).
-The volume of the thalamus compartment was further refined through three nested refinements [48] with each tetrahedron subdivided into four congruent parts, resulting in tetrahedra with diameters ranging from 0.1725 mm to 0.375 mm.
-This process led to a mesh consisting of 4.4 million nodes and 25.4 million elements (Fig.
+The quasi-static forward problem was modeled using the Complete Electrode Model (CEM) formulation [pursiainen2017forward] and solved with the Finite Element Method (FEM).
+Electrical field distributions were interpolated over the volumetric mesh following [pursiainen2016, miinalainen2019].
+All modeling, meshing, and optimization procedures were performed using the Zeffiro Interface (ZI) version 6.1.0 [pursiainen2025zeffiro], an open-source MATLAB-based pipeline that provides unstructured boundary-fitted, multi-compartment tetrahedral FE mesh generation from MRI-derived segmentations.
+2.1 Realistic Head Model
+We generated a tetrahedral-based mesh with a nominal element size of 3.0 mm from an open access T1 weighted MRI data set from a neurologically healthy adult [piastra_maria_carla_2020_3888381].
+We performed cortical segmentations using FreeSurfer [fischl2012freesurfer], supported by functions from SPM12 package [ASHBURNER20091163].
+Deep nuclei labels were extracted according to the Aseg atlas (Fig.
+1(a)) resulting in a realistic head model with 18 anatomically distinct tissue compartments.
+A double-layered cylindrical compartment with a 1.27 mm outer diameter, and a 0.5 mm thick encapsulation layer [anderson_2019_anodic] was further added and interpolated using in-house scripts.
+The cylinder functions as a proxy for setting the electrodes with their respective geometric properties.
+As our target sub-compartment would be then the anterior nucleus of the thalamus (ANT), consistent with ANT-DBS applications in epilepsy [LEHTIMAKI_2016_DBS_targetting, Jarvenpaa_2020_Improving_ANTDBS], the thalamus compartment underwent three levels of nested refinements subdividing every tetrahedron on each step into four congruent sub-elements [galazprieto_2023_mesh].
+This procedure yielded element diameters ranging from 0.1725 to 0.375 mm in the refined region and produced a final mesh comprising approximately 4.4 million nodes and 25.4 million tetrahedral elements (Fig.
 1(b)).
-We define the anterior nucleus of the thalamus (ANT) as the stimulation target, similar to ANT-DBS, a clinically established approach for treating epilepsy [8, 9].
+We assigned electrical conductivities as follows: 0.14 S/m for white matter, 0.33 S/m for gray matter, 0.0064 S/m for skull, 1.79 S/m for cerebrospinal fluid (CSF), 0.33 S/m for subcortical structures, and 0.33 S/m for skin, as reported in [dannhauer2011modeling].
+Subcortical nuclei were assumed to share electrical conductivity values of gray matter, while ventricular regions were assigned the conductivity of CSF, in accordance with [rezaei2021reconstructing] and its referenced sources.
+2.2 Lead Field Matrix
+To establish a baseline representation for spatial attenuation of the lead field and its correspondence to feasible VTA coverage, we derived an initial low-resolution lead field (LR-LF) comprising $244\times 3$ dipolar DOFs distributed throughout the thalamus.
+The spatial extent for detailed current steering was then guided by reported activation radii around the electrodes, where typical stimulation produces tissue activation on the order of approximately 2.0–4.0 mm [butson_2007_patient, miocinovic_2006_subthalamic], and may extend to roughly 3.8–4.0 mm under clinically relevant parameters [madler2012activated].
+We then establish a spherical region of 6.0 mm radius centered on the lead to represent the domain in which fine-grained modulation is physiologically meaningful.
+A high-resolution lead field (HR-LF) was subsequently generated by increasing the dipole density to $3620\times 3$ DOFs across the same anatomical region and then retaining only those sources falling within the 6.0 mm domain (Fig.
+1(c)).
+A full derivation can be found in A.1.
 (a) Segmented head model domain depicting main subcortical structures.
-(b) Refined and boundary-fitted thalamus sub-compartment.
+(b) Refined unstructured boundary-fitted thalamus sub-compartment with activity (smoothed).
 (c) Superior, coronal, and sagittal view of target domain. The interthalamic adhesion can be easily noticed.
 (d) 8- and 40-electrode contacts, respectively.
-Figure 1: (1(a)): A head model including skull, forebrain, midbrain, hindbrain structures as well as directional deep brain stimulation (dDBS) lead positioned at the right-hemispheric anterior nucleus of thalamus (ANT).
-(1(b)): Refined and boundary-fitted thalamus sub-compartment with the distribution of tetrahedral mesh edges and an interpolated current intensity distribution shown on the surface.
+Figure 1: (1(a)): A head model including skull, forebrain, midbrain, hindbrain structures as well as a directional deep brain stimulation (DBS) lead positioned at the right-hemispheric anterior nucleus of thalamus (ANT).
+(1(b)): Refined unstructured boundary-fitted thalamus sub-compartment with the distribution of tetrahedral mesh edges and an interpolated current intensity distribution shown on the surface.
 (1(c)): Placement of a 40-contact dDBS lead with a bipolar electrode configuration (anodal as red, and cathodal as blue).
-The dipolar target current (red) is depicted by the red pin, indicating the target position and orientation; a 6.0 millimeter (mm) region of interest (ROI) is displayed as an arrangement of source points (yellow squares) for High-Resolution Lead Field (HR-LF) matrix.
-(1(d)): Probing strategy of the 8- and 40-electrode contacts inspired by Abbott’s St.
-Jude Medical Infinity™ Directional Lead 6180 model and Medtronic-Sapiens, respectively.
-2.2 Lead Field Matrix
-The linear forward mapping can be formulated as a real $N\times K$ lead field matrix $\hat{\bf L}$, which relates current injections at the electrodes to the resulting volumetric current density distribution across different electrode configurations.
-This matrix characterizes the numerical relationship between the real $K\times 1$ current pattern ${\bf y=(y_{1},\ldots,y_{K})}$ and the real $N\times 1$ discretized volume current density $\hat{\bf x}=(\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{N})$.
-The latter variable represents an arrangement of uniformly distributed source points within the target domain, each with three degrees of freedom (DOFs) corresponding to three dipolar unit currents with Cartesian orientations.
-We generate a lead field matrix for two different resolution levels: (i) Low-Resolution Lead Field (LR-LF), which corresponds to a target set of 244 $\times$ 3 spatial DOFs that cover the entire thalamus compartment.
-This robust interpretation offers a functional assessment of the model, serving as a benchmark for our numerical experiments.
-To validate whether our minimal interpretation is adequate, we compare the volumetric current density distribution with a locally detailed lead field matrix, (ii) High-Resolution Lead Field (HR-LF).
-This matrix consists of 3620 $\times$ 3 DOFs distributed across the thalamus, meaning that the mutual distance between nearest neighboring DOFs is reduced by 40% compared to LR-LF.
-Due to the increased computational cost of using a larger lead field size, our examination of HR-LF is confined to a spherical region of interest with a 6.0 mm radius, co-centered with the lead model (Fig.
-1(c)).
-2.3 Probing Strategy
+The dipolar target current (red) is depicted by the red pin, indicating the target position and orientation; a 6.0 millimeter ($m$) fitting region is displayed as an arrangement of source points (yellow squares) for High-Resolution Lead Field (HR-LF) matrix.
+(1(d)): Probing strategy of the 8- and 40-electrode contacts inspired by Abbott Infinity™ and Medtronic-Sapiens, respectively.
+2.3 Optimization Framework
+The applied optimization framework aims to return a focal anodal–cathodal electrode configuration that best reproduces the volumetric current distribution associated with a prescribed target current dipole model [demunck1988, mosher2002multiple, medani_2012_ModellingDipoles].
+The framework enhances field components aligned with the target while attenuating directions dominated by nuisance activation.
+Active electrodes are selected iteratively by adjusting the current pattern $\mathbf{y}$ under a set of predefined hyperparameters and an expected number of active contacts $K$, depending on the chosen algorithm (e.g., Reciprocity Principle, L1L1 Method, or Tikhonov-regularized least squares).
+Metaheuristics are employed to explore the feasible design space and identify current patterns that best satisfy the required performance characteristics [galazprieto_2024_lattice].
+To restrict the number of admissible solutions, we impose a lower bound on the focused current density $\Gamma$ ($\Gamma\geq\Gamma_{0}=0.80~\mathrm{mA}$) to ensure sufficient activation at the target, as well as an upper bound on the nuisance current density $\Xi$ to limit excessive off-target stimulation.
+The candidate solution is thus defined as the current pattern that maximizes the field ratio between target- and nuisance-region current densities while satisfying these constraints.
+As safety constraints, the total (absolute) injected current is constrained to $\mu=4.0~\mathrm{mA}$ (6), with per-contact limits $\gamma=\pm 2.0~\mathrm{mA}$ (7).
+Electrode potentials are normalized to satisfy the net-current constraint (8).
+Lastly, the stimulation is, however, treated as time-invariant, as the objective of this work is to characterize the spatial organization of current flow and the resulting current density distributions, rather than to model pulse dynamics or frequency-dependent neuronal responses.
+For reference magnitude $\|{\bf x}_{1}\|_{2}=3.85$ A/m2 corresponds to the activation threshold reference.
+This value is obtained by dividing the average dipole moment density of pyramidal neurons (0.77 nAm/mm2) by an approximate dendritic axis length of 0.2 mm [MURAKAMI201549].
+2.3.1 Electrode Configurations
 […]
-The maximum total current injection supplied by all the active electrodes is limited to 4.0 mA (milliamperes) (5), and the maximum current per electrode contact is set to $\pm$ 2.0 mA (6).
-This can be considered a moderately typical setting, for example, 3.0 mA as a reference for monopolar patterns in [50].
-The total absolute sum of electric potential corresponding to the contacts in the lead is normalized to meet the intended maximum current injection value (7).
-As a reference threshold for neural activation, we consider the current amplitude of 3.85 A/m2 (amperes per square meter) [51], which falls within the excitation current thresholds of 2.5 to 6.0 A/m2 estimated for nerve fibers in the upper limb area of the motor cortex at frequencies of 50.0 Hz and 2.44 kHz [52].
-Our optimization framework maintains this reference threshold constant across all tested configurations.
-By holding temporal characteristics constant, we ensure that the observed variations in activation patterns arise from differences in spatial registrations and current steering, rather than from stimulation timing.
-This choice follows the framework established in previous studies by Deli et al.
-[53] and Schlaepfer et al.
-[54], where monopolar and bipolar stimulation configurations were evaluated using a consistent stimulation protocol, specifically maintaining a pulse width of 60 µs and a frequency of 130 Hz across all conditions.
-This design allowed the authors to isolate the effects attributable solely to electrode configuration by holding other key stimulation parameters constant, with amplitude as the only adjusted variable.
-2.4 Optimization Algorithms
-The aim of this study is the application of mathematical optimization algorithms to retrieve the current pattern ${\bf y}$ (4) within the local target region, aiming to optimally match the volumetric current distribution generated by the target current dipole modeling [55, 56, 51].
-To quantify the optimal electrode configurations derived from the algorithm, we compare them based on the decision variables $\Gamma$, $\Xi$, and $\Theta$, as delineated by the optimization paradigm.
-The focused current density $\Gamma$ (A/m2) is defined as
-${\Gamma}=\frac{{\bf x}_{1}^{T}{\bf L}_{1}{\bf y}}{\|{\bf x}_{1}\|_{2}}.$
-The focality of the stimulation, or field ratio, $\Theta$, is defined as the ratio of the focused current density divided by the undesired nuisance current density ($\Xi$) and is defined as
-${\Theta}=\frac{{\Gamma}}{\Xi}\quad\hbox{with}\quad\Xi=\frac{\|{\bf L}_{2}{\bf y}\|_{2}}{\sqrt{M}}.$
 2.3.2 Bipolar Configurations
-Bipolar configurations are derived through the reciprocity principle (RP) [36].
-The principle states that the maximum focused current density, i.e., $\Gamma_{\hbox{\scriptsize{max}}}$, of an applied linear system can be achieved by limiting the electrode configuration to a pair of anodal and cathodal electrode contacts (Fig.
-3).
-This configuration is obtained by multiplying the dipolar target current distribution by the transpose of the lead field matrix, selecting both the greatest positive and negative poles, and setting the remaining electrodes in the lead to zero.
-The reciprocity principle aims at maximizing current density at the expense of field ratio fidelity.
+Bipolar patterns are obtained using the reciprocity principle (RP) [FERNANDEZCORAZZA2020116403].
+The principle states that the maximal focused current density $\Gamma_{\mathrm{max}}$ for a linear quasi-static system is achieved by selecting a single anode–cathode pair.
+In practice, the target dipolar field is multiplied by the transpose of the lead field matrix, and the contacts corresponding to the strongest positive and negative entries are chosen as anode and cathode, respectively, with all remaining contacts set to zero (Fig.
+2).
+RP maximizes the delivered current density at the target but does not optimize the ratio between target activation and nuisance activation, and thus may yield suboptimal field selectivity.
 2.3.3 Multipolar Configurations
-Multipolar configurations, i.e., leads with multiple active electrode contacts, are obtained via convex optimization techniques such as L1-norm regularized L1-norm fitting (L1L1) [32] and through Tikhonov-Regularized Least Squares (TLS) [37, 38] methods.
-Both methods employ a metaheuristic search strategy to explore only a carefully chosen fraction of an otherwise huge search space—trading guaranteed optimality for scalability and adaptability—to handle large, ill-behaved, or poorly understood problems within practical time and memory constraints [57].
-The candidate solution is the one that can achieve the highest ratio between the (target) current density at the stimulation site and that in the surrounding adjacent areas (referred here to as nuisance current).
-The determination of this solution is defined by tuning a set of predefined hyperparameters along with the current pattern ${\bf y}$, subject to specific metacriteria, namely, maximizing the field ratio while ensuring $\Gamma\geq\Gamma_{0}>0$, thereby maintaining sufficient intensity at the target location (with $\Gamma_{0}=0.80$ mA in this case).
-Without a lower bound for intensity $\Gamma$, the intensity of the $\Theta$-maximizer is likely to diminish.
-A rough upper bound is provided as a L1-norm fitting threshold to suppress nuisance current density $\Xi$.
-In contrast, traditional least-squares formulations tend to produce smoothly distributed currents, which may be clinically suboptimal, as focal and sparse activation is often preferred to minimize side effects and enhance therapeutic precision.
-Moreover, least-squares approaches are sensitive to noise and outliers, often yielding non-robust solutions [58].
-Detailed formulation of both the L1L1 and TLS methods is provided in A.
+Multipolar configurations are derived using the L1-norm regularized L1-norm fitting (L1L1) method [galazprieto_2022_L1vsL2] and Tikhonov-regularized least squares (TLS) [dmochowski2011optimized].
+Both methods assume a predetermined set of candidate active contacts and estimate the corresponding stimulation amplitudes under polarity, charge, and safety constraints (detailed in A.2).
+The L1L1 formulation imposes an L1 penalty on both the stimulation vector and the residual, which enhances robustness to lead-field uncertainty and promotes sparse, interpretable support.
+TLS, in contrast, employs an L2 penalty that stabilizes the solution in the presence of correlated columns and numerical ill-conditioning.
+Because the full combinatorial search space of possible contact subsets grows exponentially with $K$, both formulations employ a metaheuristic support-selection strategy to efficiently explore only a tractable subset of candidate configurations while maintaining practical computational cost [galazprieto_2024_lattice].
+8-contact Electrode Configuration
+Refer to caption RPRefer to captionRefer to caption Refer to caption L1L1(A)Refer to captionRefer to caption Refer to caption L1L1(B)Refer to captionRefer to caption Refer to caption TLSRefer to captionRefer to caption
+40-contact Electrode Configuration
+Refer to caption RPRefer to captionRefer to caption Refer to caption L1L1(A)Refer to captionRefer to caption Refer to caption L1L1(B)Refer to captionRefer to caption Refer to caption TLSRefer to captionRefer to caption
+Figure 2: Examples of current injection patterns obtained using the Reciprocity Principle (RP), metaheuristic L1-norm regularized L1-norm fitting (L1L1) method under two different nuisance current thresholds (L1L1(A) with $\varepsilon=\{0:-160\}$ and L1L1(B) with $\varepsilon=\{0:-10\}$), and the Tikhonov-Regularized Least Square (TLS) method.
+The target position and orientation (parallel to the lead) is indicated by the red upward pin.
+The bar plots display the resulting electrode current amplitudes, while the contact maps visualize their spatial distribution over the electrodes; cathodal (blue), anodal (red), and inactive (gray/green).
+The RP method yields a bipolar configuration maximizing focused current density $\Gamma$, while the remaining methods searches for a multipolar configuration maximizing focality $\Theta=\Gamma/\Xi$ under an a priori constraint $\varepsilon$ on the nuisance current density $\Xi$, determined by the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ and the lead field attenuation.
+The L1L1(B) result reflects more strongly constrained $\Xi$ (smaller dynamic range), activating fewer contacts whereas L1L1(A) and TLS allow higher $\Theta$ values and broader excitation patterns.
+The voltage/current on the electrode contacts are $\pm 2.0$ mA (milliamperes).
 2.4 Uncertainty Modeling
-Uncertainty refers to various physiological and technical factors that directly or indirectly affect the accuracy and reliability of the stimulation, including imprecision in lead placement, patient-specific anatomical variation, and inhomogeneities in tissue conductivity.
-In this study, all sources of uncertainty affecting the optimization problem are modeled as additive perturbations to the lead field.
-While this Gaussian noise model is a simplification, it enables a tractable first-step analysis of robustness.
+Uncertainty refers to various physiological and technical factors that affect the accuracy and reliability of the stimulation, including imprecision in lead placement, patient-specific anatomical variation, and inhomogeneities in tissue conductivity [schmidt2016uncertainty, athawale2019statistical].
+Potential sources affecting the optimization framework are modeled as additive perturbations to the lead field matrix through Gaussian noise model.
+While this is a simplification, it enables a tractable first-step analysis of robustness.
 We acknowledge that in practice, the true uncertainty arises from a variety of unrelated sources.
+We postulate that the volumetric current density distribution can be optimized for DOFs indexed by $i$ whose corresponding reduced lead field row ${\bf\tilde{l}}_{i}$ has an L2-norm above a threshold.
+This row is first reduced and zero-averaged, taking the form
+$\tilde{\bf l}_{i}=\left(L_{i,1}-\overline{L}_{i},L_{i,2}-\overline{L}_{i},\dots,L_{i,K}-\overline{L}_{i}\right),$
+where $\overline{L}_{i}=K^{-1}\sum_{j=1}^{K}L_{i,j}$ is the mean of the row entries.
+The threshold is determined via a peak signal-to-noise ratio (PSNR) constraint, ensuring that only DOFs contributing significantly to the lead field are included in the optimization.
+The feasible set of DOFs is defined as
+$\mathcal{V}_{\delta}=\{i\,|\,\|{\bf\tilde{l}}_{i}\|_{2}\geq\delta\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}\},$
+where $\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}$ is the largest L2-norm among all the zero-averaged rows in the lead field rows.
+The parameter $\delta$ sets a relative threshold, filtering out DOFs whose contributions are too weak to be meaningful.
+When $\delta$ is interpreted as the activation threshold, $\mathcal{V}_{\delta}$ approximates the maximum VTA (Fig.
+3(a)).
+Because lead field attenuation and uncertainty impose limits on stimulation accuracy, the PSNR constraint effectively defines the theoretical performance limit of the optimization.
+If the target site corresponds to index $\ell$, the maximum achievable dynamic range $\varepsilon^{-1}$ between the current density at the target and surrounding regions is approximately
+$\varepsilon\geq\delta\,\|{\bf\tilde{l}}_{\ell}\|_{2}^{-1}\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}.$
+This indicates that the minimum contrast between the target and non-target regions is constrained by both $\delta$ and the structure of the lead field matrix (Fig. 3(b)).
 […]
@@
 (b) The set $\mathcal{V}_{\delta}$ illustrated for the volumetric model of this study with respect to PSNR $\delta^{-1}$ levels from 10 to 40 dB.
-Figure 2: (2(a)): A schematic of how the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ relates to the set $\mathcal{V}_{\delta}$ and to the approximate maximum dynamic range $\varepsilon^{-1}$ of the volumetric stimulation current field.
-(2(b)): The set $\mathcal{V}_{\delta}$ shown for PSNR 10–40 dB, providing a rough approximation of the maximum feasible coverage of volume of tissue activated (VTA) for an activation threshold set by $\delta$.
-8-contact Electrode Configuration
-Refer to caption RPRefer to captionRefer to caption Refer to caption L1L1(A)Refer to captionRefer to caption Refer to caption L1L1(B)Refer to captionRefer to caption Refer to caption TLSRefer to captionRefer to caption
-40-contact Electrode Configuration
-Refer to caption RPRefer to captionRefer to caption Refer to caption L1L1(A)Refer to captionRefer to caption Refer to caption L1L1(B)Refer to captionRefer to caption Refer to caption TLSRefer to captionRefer to caption
-Figure 3: Examples of current injection patterns obtained using the Reciprocity Principle (RP), metaheuristic L1-norm regularized L1-norm fitting (L1L1) method under two different nuisance current thresholds (L1L1(A) with $\varepsilon=\{0:-160\}$ and L1L1(B) with $\varepsilon=\{0:-10\}$), and the Tikhonov-Regularized Least Square (TLS) method.
-The target position and orientation (parallel to the lead) is indicated by the red upward pin.
-The bar plots display the resulting electrode current amplitudes, while the contact maps visualize their spatial distribution over the electrodes; cathodal (blue), anodal (red), and inactive (gray/green).
-The RP method yields a bipolar configuration maximizing focused current density $\Gamma$, while the remaining methods searches for a multipolar configuration maximizing focality $\Theta=\Gamma/\Xi$ under an a priori constraint $\varepsilon$ on the nuisance current density $\Xi$, determined by the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ and the lead field attenuation.
-The L1L1(B) result reflects more strongly constrained $\Xi$ (smaller dynamic range), activating fewer contacts whereas L1L1(A) and TLS allow higher $\Theta$ values and broader excitation patterns.
-The voltage/current on the electrode contacts are ±$\pm 2.0$ mA (milliamperes).
-We postulate that the volumetric current density distribution can be optimized for those degrees of freedom indexed by $i$, where the L2-norm (2-norm) $\|{\bf\tilde{l}}_{i}\|_{2}$ of the corresponding reduced lead field matrix row remains above the threshold.
-This row, denoted as ${\bf l}_{i}$, is first reduced (as described in Appendix 9) and then zero-averaged, taking the form
-$\tilde{\bf l}_{i}=\left(L_{i,1}-\overline{L}_{i},L_{i,2}-\overline{L}_{i},\dots,L_{i,K}-\overline{L}_{i}\right),$
-where $\overline{L}_{i}=K^{-1}\sum_{j=1}^{K}L_{i,j}$ represents the mean of the row entries.
-The threshold for optimization is determined by a peak signal-to-noise ratio (PSNR) constraint, ensuring that only degrees of freedom with a sufficiently strong contribution to the lead field are considered in the optimization process.
-We define the feasible set of degrees of freedom as
-$\mathcal{V}_{\delta}=\{i\,|\,\|{\bf\tilde{l}}_{i}\|_{2}\geq\delta\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}\},$
-where $\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}$ is the maximum L2-norm of the zero-averaged lead field rows.
-The parameter $\delta$ sets a relative threshold, filtering out degrees of freedom whose contributions are too weak to be considered meaningful.
-If $\delta$ is chosen to represent neural activation threshold, then $\mathcal{V}_{\delta}$ serves as an approximation of the maximum VTA (Fig.
-2(a)).
-Since lead field attenuation and uncertainty impose fundamental limits on stimulation accuracy, the PSNR constraint defines the theoretical limit of the optimization performance.
-If the target stimulation site corresponds to the $\ell$-th index, then the maximum achievable dynamic range $\varepsilon^{-1}$ between the current density at the target site and its surroundings is approximately
-$\varepsilon\geq\delta\,\|{\bf\tilde{l}}_{\ell}\|_{2}^{-1}\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}.$
-This relationship indicates that the minimum contrast in current density between the target location and surrounding regions is constrained by $\delta$ and the structure of the lead field matrix (Fig.
-2(b)).
+Figure 3: (3(a)): A schematic of how the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ relates to the set $\mathcal{V}_{\delta}$ and to the approximate maximum dynamic range $\varepsilon^{-1}$ of the volumetric stimulation current field.
+(3(b)): The set $\mathcal{V}_{\delta}$ shown for PSNR 10–40 dB, providing a rough approximation of the maximum feasible coverage of volume of tissue activated (VTA) for an activation threshold set by $\delta$.
+2.5 Hardware
+Dell Precision 5820 Workstation equipped with 256 GB of RAM, a 10-core Intel i9–10900X CPU, and an NVidia Quadro RTX 4000 GPU. The GPU were used to accelerate both forward and inverse calculations.
 2.6 Numerical Experiments
-Our numerical experiments are designed to evaluate the performance of the proposed optimization methods with respect to: (i) simulated 8-contact and 40-contact DBS lead configurations, (ii) noiseless and noisy forward mappings, (iii) orientation of the current dipole in the model (parallel or perpendicular with respect to the lead array) [17], and (iv) algorithm efficiency.
-The hardware used for these experiments includes a Dell Precision 5820 Workstation with 256 GB of RAM, a 10-core Intel i9–10900X CPU, and an NVidia Quadro RTX 4000 GPU.
-The GPU (graphics processing unit) was utilized to accelerate both forward and inverse computations.
+Our numerical experiments evaluate the proposed optimization methods with respect to: (i) 8- and 40-contact DBS lead configurations, (ii) noiseless and noisy forward mappings, (iii) dipole orientation relative to the lead array (parallel or perpendicular) [Janson_2020_Activation], and (iv) algorithm efficiency.
 Table 1: Summary of the proposed optimization methods.
-While the RP provides a bipolar reciprocal of the target, both L1L1 and TLS constitute a metaheuristic search over a two-dimensional hyperparameter lattice.
-L1L1 is subdivided into L1L1(A) and L1L1(B) based on two alternative a priori assumptions of the feasible $\varepsilon$ threshold range which is [-160, 0] and [-10, 0] dB, respectively.
-The regularization and the L2-norm weighting parameter intervals are chosen as in [32].
-CPU time refers to the computing time needed to evaluate a single optimizer with fixed hyperparameter values for HR-LF with the 40-contact lead configuration.
+RP provides a bipolar reciprocal of the target, whereas L1L1 and TLS perform an extensive metaheuristic search.
+L1L1 method is subdivided into L1L1(A) and L1L1(B) according to two alternative a priori assumptions for the feasible $\varepsilon$ threshold range: [-160, 0] dB and [-10, 0] dB, respectively.
+The regularization and L2-norm weighting parameter ranges are adopted from [galazprieto_2022_L1vsL2].
+CPU time refers to the computation time needed to evaluate a single optimizer with fixed hyperparameters for the HR-LF using the 40-contact lead.
 Hyperparameter CPU
@@
 $\beta\in[-50,40]$ Weight
-The L1L1 method is examined using two alternative a priori assumptions for the feasible $\varepsilon$ threshold range: [-160, 0] dB and [-10, 0] dB, respectively.
-We refer to these settings as ”L1L1(A)” and ”L1L1(B)”.
-The former setting neglects the effect of uncertainty on the field ratio $\Theta$, while the second assumes that uncertainty is included within 10 dB dynamic range.
-As for TLS method, the regularization parameter $\gamma$ and the L2-norm weighting parameter $\beta$ intervals are chosen as in [32] (Table 1).
-Gaussian noise is added to the lead field to examine spatial sensitivity; the effect of noise is assessed in an overall setting, where the candidate solution is determined for each target DOF of LR-LF under PSNR 40.0 dB noise corruption.
-While anatomical variations may cause spatially correlated errors, the implementation of noise offers a generic and tractable way to assess noise-robustness when the exact structure of uncertainty is unknown.
-It serves as a standard baseline for evaluating sensitivity to model perturbations.
-The noise sensitivity is examined for two specific target positions (I) and (II) (Fig.
-4) using PSNRs of 40.0 and 50.0 dB, along the HR-LF.
-When interpreted in terms of $\mathcal{V}_{\delta}$ (Figure 2(b)), PSNR 40.0 dB (50.0 dB) corresponds to the assumption that each possible target position in the set $\mathcal{V}_{-30\text{ dB}}$ ($\mathcal{V}_{-40\text{ dB}}$) can be stimulated with at least a 10 dB dynamic range $\varepsilon^{-1}$ for $\Theta$.
-This assumption aligns well with the $\varepsilon$ range of $[-10,0]$ selected for L1L1(B).
+The L1L1 method is tested under two alternative a priori assumptions for the feasible $\varepsilon$ threshold range: [-160, 0] dB (L1L1(A)) and [-10, 0] dB (L1L1(B)).
+The first neglects uncertainty in the field ratio $\Theta$, while the second accounts for variability within a 10 dB dynamic range.
+TLS uses regularization parameter $\gamma$ and L2-norm weighting parameter $\beta$ intervals as in [galazprieto_2022_L1vsL2] (Table 1).
+Detailed formulation of both the L1L1 and TLS methods is provided in A.2 and A.3, respectively.
+To examine sensitivity to measurement and modeling errors, Gaussian noise is added to the lead field.
+Candidate solutions are computed for each target DOF of the LR-LF under PSNR 40.0 dB corruption.
+While anatomical variations can introduce spatially correlated errors, this additive noise provides a tractable baseline to assess robustness to model perturbations.
+Noise sensitivity is specifically analyzed for two target positions (I) and (II) (Fig.
+4) along the HR-LF using PSNRs of 40.0 and 50.0 dB.
+Interpreted in terms of $\mathcal{V}_{\delta}$, PSNR 40.0 dB (50.0 dB) corresponds to the assumption that each target in $\mathcal{V}_{-30\text{ dB}}$ ($\mathcal{V}_{-40\text{ dB}}$) can be stimulated with at least a 10 dB dynamic range $\varepsilon^{-1}$ for $\Theta$, which aligns with the $\varepsilon$ range of [-10, 0] used for L1L1(B).
 Figure 4: Visualization of the two synthetic stimulation targets, Target (I) and Target (II), analyzed separately in numerical experiments.
@@
 Bilateral thalamic structures are shown in gray for anatomical reference.
+The distances are shown with respect to the boundary cylindrical probe illustrated by the ring around the center point.
 LR-LF HR-LF
@@
 Figure 5: Volumetric optimization results for the 8-contact probe and noiseless lead fields.
-Decision variable $\Gamma$, $\Xi$ and $\Theta$ for optimizers found via RP, TLS, L1L1(A), and L1L1(B) method illustrated as a function of the target position and orientation (parallel or perpendicular w.r.t.
+Decision variable $\Gamma$, $\Xi$ and $\Theta$ for optimizers found via Reciprocity Principle (RP), L1-norm regularized L1-norm fitting (L1L1) evaluated as L1L1(A) and L1L1(B) with feasible $\varepsilon$ threshold range: [-160, 0] dB and [-10, 0] dB, respectively, and Tikhonov Regularized Least Squares (TLS) methods illustrated as a function of the target position and orientation (parallel or perpendicular w.r.t.
 the probe).
-Left: distributions for a (i) Low-Resolution Lead Field (LR-LF) where the spatial degrees of freedom (DOFs) cover both thalamus lobes.
-Right: distribution for a (ii) High-Resolution Lead Field (HR-LF) where the DOFs is confined to a region of interest within a radius of 6.0 mm (millimeters) from the center of the lead.
+Left: distributions for a Low-Resolution Lead Field (LR-LF) where the spatial degrees of freedom (DOFs) cover both thalamus lobes.
+Right: distribution for a High-Resolution Lead Field (HR-LF) where the DOFs is confined to a region of interest within a radius of 6.0 mm (millimeters) from the center of the lead.
 The optimal electrode configuration found by the L1L1 method maximizes the field ratio $\Theta$ subject to $\Gamma>=\Gamma_{0}$.
@@
 6) leads.
-The direction of the dipole current plays a key role in shaping electric field targeting.
+The direction of the current dipole plays the role in shaping electric field targeting.
 The anode–cathode pair exhibiting the strongest positive and negative values may activate electrodes located as close as $0.189\pm 0.009\,\text{mm}$ or as far as $6.375\pm 0.375\,\text{mm}$ apart across the lead.
-This effect becomes evident when comparing the corresponding distributions: in the parallel orientation, the targeting is highly concentrated near the tip of the probe, whereas the perpendicular orientation results in a broader spatial spread, with activity shifted medially relative to the lead.
-The bipolar configurations obtained via RP yield the overall highest focused current density $\Gamma$; however, this comes at the expense of focality, as indicated by elevated nuisance current $\Xi$ and a reduced field ratio $\Theta$.
+With the parallel orientation, the targeting is highly concentrated near the tip of the probe, whereas the perpendicular orientation results in a broader spatial spread, with activity shifted medially relative to the lead.
+Bipolar configurations obtained via RP yield the highest focused current density $\Gamma$; however, this comes at the expense of focality, as indicated by elevated nuisance current $\Xi$ and a reduced field ratio $\Theta$.
 This trade-off is understandable, as the multipolar configurations are designed to maximize the field ratio $\Theta$ subject to the constraint $\Gamma\geq\Gamma_{0}>0$, with $\Gamma_{0}=0.8\,\text{mA}$, thereby allowing activation of more than two contacts.
-In the noiseless setting, the L1L1(B) reconstructions with $\varepsilon\in[-10,0]\,\text{dB}$ yield slightly higher values of $\Gamma$ and $\Xi$, but exhibit a more moderate $\Theta$ compared to L1L1(A) with $\varepsilon\in[-160,0]\,\text{dB}$.
-Nonetheless, both L1L1(A) and L1L1(B) demonstrate superior nuisance current suppression relative to TLS and RP.
-Figure 7 and Table 2 compare the noiseless and noisy ($\text{PSNR}=40\,\text{dB}$) results for $\Gamma$ and $\Theta$ obtained using L1L1(A) and L1L1(B) and LR-LF.
-The results suggest that, unlike L1L1(A), L1L1(B) is more robust to lead field noise, yielding similar outcomes in both settings.
-In contrast, L1L1(A), which uses a smaller lower bound for $\varepsilon$, tends to overfit the noiseless data.
-This leads to unrealistically high values of the field ratio $\Theta$: up to $27.07$ in the 8-contact configuration and as high as $536.05$ in the 40-contact case.
-However, in the presence of noise, the results produced by L1L1(A) fall within a similar range to those of L1L1(B), indicating that the overfitting is specific to the noiseless scenario.
-Figure 8 further examines the noise sensitivity of RP, TLS, L1L1(A), and L1L1(B) for two specific target positions, illustrated in Figure 4, covering a sample of twenty different noise realizations for both PSNR 40 and 50 dB.
-The results demonstrate that in a noisy situation, L1L1(A) and L1L1(B) provide similar distributions for $\Gamma$ and $\Theta$, hence, also for $\Xi$.
-Compared to RP and TLS, the L1L1 approach seems to provide a somewhat elevated field ratio $\Theta$ with a difference that is only marginal or absent with PSNR 40 dB, but observable with PSNR 50 dB.
+Figure 7 and Table 2 compare noiseless and noisy ($\text{PSNR}=40\,\text{dB}$) results for $\Gamma$ and $\Theta$ using the L1L1 method over a LR-LF.
+In the noiseless setting, the L1L1(B) reconstructions with $\varepsilon\in[-10,0]\,\text{dB}$ yield slightly higher $\Gamma$ and $\Xi$ values, and moderate $\Theta$ compared to L1L1(A) with $\varepsilon\in[-160,0]\,\text{dB}$.
+L1L1(A) provided unrealistically high values of the field ratio $\Theta$: up to $27.07$ in the 8-contact directional DBS leads and as high as $536.05$ in the 40-contact case indicating a case of overfitting.
+Figure 8 further examines the noise sensitivity of the applied optimization methods using two specific target positions, illustrated in Figure 4, covering a sample of twenty different noise realizations for both PSNR 40 and 50 dB.
+The results demonstrate that in a noisy situation, L1L1(A) and L1L1(B) provide similar distributions for $\Gamma$ and $\Theta$.
+Compared to RP and TLS, the L1L1 approach seems to provide elevated field ratio $\Theta$ with a difference that is only marginal or absent with PSNR 40 dB, but observable with PSNR 50 dB.
 Accordingly, the potential benefit of using L1L1 depends critically on the level of modeling uncertainty, and it vanishes when this uncertainty becomes excessive.
 Of the four optimization approaches tested, L1L1(B) provides the closest agreement between noiseless and noisy outcomes and proves particularly effective in maximizing the field ratio $\Theta$ in the presence of noise.
-Table 2: The maximum values of the decision variable (DV) $\Theta$ (field ratio) found by L1L1(A) and L1L1(B) searches and the corresponding values of $\Gamma$ (focused current density) for the low-resolution lead field (LR-LF) and its target set (Figure 7) covering both parallel and perpendicular target orientations of the DBS lead in both noiseless and noisy (PSNR 40 dB) setting.
+Table 2: Maximum focused current density and field ratio described as $\Gamma_{max}$ and $\Theta_{max}$, respectively, for a 8- and 40-contacts multipolar DBS lead using the L1L1 method over a low-resolution lead field (LR-LF), with current dipole targeting aligned parallel and perpendicular to the DBS lead, for both noiseless (No Noise) and noisy (PSNR 40 dB) settings.
 8-contacts Multipolar
@@
 DV Orientation No Noise PSNR 40 dB No Noise PSNR 40 dB
-$\Gamma$ Parallel 2.42 2.27 2.27 2.27
+$\Gamma_{max}$ Parallel 2.42 2.27 2.27 2.27
 Perpendicular 3.04 2.35 2.28 2.34
-$\Theta$ Parallel 21.53 5.56 13.08 5.48
+$\Theta_{max}$ Parallel 21.53 5.56 13.08 5.48
 Perpendicular 27.07 7.49 13.18 7.22
@@
 DV Orientation No Noise PSNR 40 dB No Noise PSNR 40 dB
-$\Gamma$ Parallel 2.20 1.74 2.02 2.24
+$\Gamma_{max}$ Parallel 2.20 1.74 2.02 2.24
 Perpendicular 2.24 2.29 3.37 2.23
-$\Theta$ Parallel 526.44 7.56 9.17 8.49
+$\Theta_{max}$ Parallel 526.44 7.56 9.17 8.49
 Perpendicular 536.05 9.96 10.97 8.47
 8-contact Electrode Configuration
-  Noiseless lead field Refer to caption A1.
-Focus $\Gamma$ (A/m2)     Refer to caption B1.
-Ratio $\Theta$ (rel.) Noisy lead field Refer to caption C1.
-Focus $\Gamma$ (A/m2)     Refer to caption D1.
-Ratio $\Theta$ (rel.)
+Noiseless lead field
+A1. Focus $\Gamma$ (A/m2)
+B1. Ratio $\Theta$ (rel.)
+Noisy lead field
+C1. Focus $\Gamma$ (A/m2)
+D1. Ratio $\Theta$ (rel.)
 40-contact Electrode Configuration
-  Noiseless lead field Refer to caption A2.
-Focus $\Gamma$ (A/m2)     Refer to caption B2.
-Ratio $\Theta$ (rel.) Noisy lead field Refer to caption C2.
-Focus $\Gamma$ (A/m2)     Refer to caption D2.
-Ratio $\Theta$ (rel.)
+Noiseless lead field
+A2. Focus $\Gamma$ (A/m2)
+B2. Ratio $\Theta$ (rel.)
+Noisy lead field
+C2. Focus $\Gamma$ (A/m2)
+D2. Ratio $\Theta$ (rel.)
 L1L1(A), $\varepsilon\in[-160,0]$, Parallel
@@
 8-contact Electrode Configuration
-  Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (I) Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (II)
+Parallel
+Focus $\Gamma$ (A/m2)
+Ratio $\Theta$ (rel.)
+Perpendicular
+Focus $\Gamma$ (A/m2)
+Ratio $\Theta$ (rel.)
+Target (I)
+Parallel
+Focus $\Gamma$ (A/m2)
+Ratio $\Theta$ (rel.)
+Perpendicular
+Focus $\Gamma$ (A/m2)
+Ratio $\Theta$ (rel.)
+Target (II)
 40-contact Electrode Configuration
-  Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (I) Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (II)
+Parallel
+Focus $\Gamma$ (A/m2)
+Ratio $\Theta$ (rel.)
+Perpendicular
+Focus $\Gamma$ (A/m2)
+Ratio $\Theta$ (rel.)
+Target (I)
+Parallel
+Focus $\Gamma$ (A/m2)
+Ratio $\Theta$ (rel.)
+Perpendicular
+Focus $\Gamma$ (A/m2)
+Ratio $\Theta$ (rel.)
+Target (II)
 Figure 8: Boxplots showing the focused current density $\Gamma$ and field ratio $\Theta$ obtained with a noise-corrupted HR-LF and the 8-contact electrode configuration for test targets (I) and (II).
-The optimizers have been found through RP, TLS, L1L1(A), and L1L1(B) search.
+The optimizers have been found through RP, L1L1(A), L1L1(B), TLS methods.
 Each bar collates a sample of twenty decision variable values corresponding to optimizers obtained with as many independent lead field noise realizations (PSNR 40 or 50 dB).
 The box shows the interquartile range (IQR) between the first quantile $Q1$ and the third one $Q3$, with a horizontal line at the median.
-The whiskers extend from the edges of the box to the smallest and largest values within $[Q1-1.5\times\text{IQR},Q3+1.5\times\text{IQR}]$, where $\text{IQR}=Q3-Q1$.
+The whiskers extend from the edges of the box to the smallest and largest values within $Q1-1.5\times\text{IQR},Q3+1.5\times\text{IQR}]$, where $\text{IQR}=Q3-Q1$.
 Data points outside this range are considered outliers and are plotted individually as markers.
@@
 4 Discussion
-This study explores the application of the recently introduced metaheuristic L1L1 method [32], to optimize the current patterns for DBS leads.
-The core focus lies in L1L1’s capacity to incorporate a priori uncertainty through an explicit constraint, offering a notable advantage over traditional approaches, which often fail to account for such variability [37, 38, 36, 59].
-Specifically, we define the uncertainty through a threshold on the nuisance current density, which is determined by the relative attenuation of the lead field with respect to a target site.
-This attenuation is quantified in decibels and corresponds to typical falloff patterns observed in VTA.
-By reducing the dependence on conventional parameter tuning methods, such as trial-and-error [60, 61], this approach paves the way for more sophisticated techniques to make more informed decisions, even in the presence of inherent uncertainties in DBS settings and simulations [62, 63, 64, 65].
-Optimization was carried out using three parameters: the focused current density $\Gamma$, the nuisance current density $\Xi$, and field ratio $\Theta=\Gamma/\Xi$.
-To assess robustness and focality under uncertainty, the recently introduced L1L1 method was applied in two variants, L1L1(A) and L1L1(B), differing in the range of the nuisance current threshold $\varepsilon$, which reflects varying levels of a priori modeling uncertainty.
-Specifically, L1L1(B) constrains $\varepsilon$ to the interval $[-10,0]$ dB, accounting for higher uncertainty, whereas L1L1(A) utilizes a broader range $[-160,0]$ dB.
-For comparison, the Reciprocity Principle (RP) and Tikhonov-Regularized Least Squares (TLS) methods were included as reference approaches.
-The results reveal that the spatial distributions of $\Gamma$, $\Xi$, and $\Theta$ are strongly influenced by both the optimization method and the orientation of the dipole current.
-Parallel configurations resulted in highly localized activation near the tip of the lead, whereas perpendicular configurations yielded a broader, medially shifted targeting.
-A comparative analysis across varying noise levels demonstrated that L1L1 methods consistently outperformed RP and TLS in maintaining higher and more stable values of $\Theta$, especially at higher PSNR values.
-In noiseless conditions, solutions obtained with a low nuisance threshold $\varepsilon\in[-160,0]\,\text{dB}$, as in L1L1(A), often exhibited unrealistically high field ratio values $\Theta$, sometimes up to fifty times greater than those derived using the more conservative threshold $\varepsilon\in[-10,0]\,\text{dB}$ in L1L1(B).
-This tendency suggests overfitting when the model is not constrained by uncertainty.
-Introducing this noise regularized the solution space, preventing the algorithm from converging to over-idealized solutions that mirrored the original model structure.
-These findings underscore the significance of explicitly modeling uncertainty and justify the constraint imposed on the nuisance current in L1L1(B), which leads to more realistic and generalizable solutions.
-Overall, these observations highlight the potential of L1L1(B) as a robust and uncertainty-aware optimization strategy for dDBS lead configuration.
-The estimates for the set $\mathcal{V}_{\delta}$, where the lead field amplitude attenuates by a factor $\delta$ for $\delta\in[-40,-10]$ dB, align closely with established knowledge regarding VTA.
-Specifically, for $\delta\approx-40\,\text{dB}$, our model predicts attenuation within a spatial region consistent with a VTA radius of approximately 4.0 mm, in line with the findings of Butson’s [33] and McIntyre’s [20].
-This agreement suggests that the noise levels used, with PSNR values of 40 dB and 50 dB, represent reasonable bounds for physiological uncertainty.
-At 40 dB, the dynamic range of $\Theta$ is determined by the maximum lead field attenuation within the VTA, while 50 dB further supports the assumption that each position within the VTA can be independently and coherently targeted.
-In this context, the L1L1 method, particularly L1L1(B), demonstrated superior targeting capabilities compared to RP and TLS, especially under noisy conditions.
-When comparing the electrode configurations, increasing the number of contact points generally enhances both adaptability and control over the resulting electric field.
-This improvement is driven by several factors, including reduced distances between electrodes and target locations within the lead field matrix, better alignment of the induced electric field with the target current direction, and a greater capacity for flexible fitting due to the increased number of degrees of freedom.
-The findings of this study provide a proof-of-concept for integrating a priori uncertainty modeling into numerically simulated dDBS optimization.
-When relevant sources of uncertainty, such as anatomical variability, conductivity inhomogeneities, or lead placement deviations, are known or can be estimated, the L1L1 method might be employed to improve the focality and directionality of stimulation, enabling more selective targeting of desired structures while reducing activation in surrounding regions.
-In clinical practice, such improvements might lead to more personalized and effective treatment strategies, enhancing therapeutic precision, and minimizing side effects [10].
+This study explores the application of the recently introduced metaheuristic L1L1 method [galazprieto_2022_L1vsL2] to optimize current patterns and electrode configurations for DBS leads.
+The L1L1 method was applied in two variants, L1L1(A) constrains $\varepsilon$ to the interval $160,0]\,\text{dB}$ and L1L1(B) to $10,0]\,\text{dB}$ reflecting different levels of a priori modeling uncertainty [schmidt2016uncertainty, athawale2019statistical].
+For comparison, optimization performance using Reciprocity Principle (RP)  [FERNANDEZCORAZZA2020116403] and Tikhonov Regularized Least Squares (TLS) [dmochowski2011optimized] methods were included.
+The core advantage in applying the L1L1 method is the capacity to incorporate a priori uncertainty through an explicit constraint, offering an advantage over traditional approaches, such as the least-square methods, which often do not account for such variability.
+Within the optimization framework, the focused current density $\Gamma$, nuisance current density $\Xi$, and field ratio $\Theta=\Gamma/\Xi$ are implemented as the three metrics to measure spatial current distributions influenced by the quality of the forward modelling, position and orientation of the current dipole, and uncertainness which is defined through a threshold on the nuisance current density, determined by the relative attenuation of the lead field matrix with respect to a target stimulation site.
+This attenuation is quantified in decibels and corresponds to typical falloff patterns observed in the volume of tissue activated (VTA) [zhang2020steering].
+In noiseless conditions, solutions using a more permissive threshold value (L1L1(A) case) often exhibited unrealistically high field ratio values, approximately fifty times greater than those derived using a more conservative threshold (L1L1(B) case).
+The addition of noise promoted the optimizer’s reliance from convergence towards over-idealized solutions, forcing the algorithm to select current patterns that remain effective even under slight model variations, producing more physiologically realistic distributions.
+The estimates for the set $\mathcal{V}_{\delta}$, where the lead field amplitude attenuates by a factor $\delta$ for $\delta\in[-40,-10]\,\text{dB}$, align closely with established knowledge regarding VTA.
+For $\delta\approx-40\,\text{dB}$, the model predicts attenuation within a spatial region consistent with a VTA radius of approximately 4.0 mm, in line with Butson and McIntyre [butson_2007_patient], Miocinovic et al.
+[miocinovic_2006_subthalamic], and Mädler and Coenen [madler2012activated].
+This suggests that the noise levels used (PSNR 40 dB and 50 dB) represent reasonable bounds for physiological uncertainty.
+At PSNR 40 dB, the dynamic range of $\Theta$ is determined by the maximum lead field attenuation within the VTA, whereas PSNR 50 dB supports the assumption that each position within the VTA can be independently and coherently targeted.
+When comparing electrode configurations, increasing the number of contact points generally improves adaptability and control over the resulting electric field.
+Parallel configurations, i.e., alignment of the current dipole with the DBS lead, produced localized activation near the tip of the lead, whereas perpendicular configurations yielded a broader medially shifted targeting.
+This is driven by reduced distances between electrodes and target locations, better alignment of the induced field with the target current direction, and greater flexibility due to more DOFs.
+Overall, the findings provide a proof-of-concept for integrating a priori uncertainty modeling into numerically simulated DBS optimization.
+When sources of uncertainty, such as anatomical variability, conductivity inhomogeneities, or lead placement deviations, are known or estimated, the L1L1 method can improve focality and directionality of stimulation, enabling more selective targeting of desired structures while minimizing off-target activation.
+Such improvements have the potential to enhance therapeutic precision and reduce side effects in clinical practice [fasano2021experience].
 4.1 Limitations and Future Work
@@
 First, the threshold values used to define acceptable nuisance current levels are heuristic and may not fully capture the complexity of individual patient anatomies or tissue conductivities.
-This simplification assumes spatial uniformity in attenuation, which may not hold, for example, in the presence of anisotropic or heterogeneous tissue properties [44, 66].
-Additionally, the approach relies on a static surrogate model of the VTA and does not directly incorporate a probabilistic model of uncertainty [62], which would allow for a more comprehensive treatment of variability.
-Potential alternative surrogate approaches exist, such as, threshold-based curve fitting [67], artificial neural networks [68], and statistical emulators such as Gaussian process classifiers [69, 70], which approach VTA in different ways.
-Compared to the reference methods of this study, the L1L1 method is also slightly more demanding in terms of computing time.
+This simplification assumes spatial uniformity in attenuation, which may not hold, for example, in the presence of anisotropic or heterogeneous tissue properties [dannhauer2011modeling].
+Additionally, the approach relies on a static surrogate model of the VTA and does not directly incorporate a probabilistic model of uncertainty [schmidt2016uncertainty, athawale2019statistical], which would allow for a more comprehensive treatment of variability.
+Potential alternative surrogate approaches exist, such as, threshold-based curve fitting [butson_2006_electrodesign], and statistical emulators such as Gaussian process classifiers [DeLaPava_2017_accelerating, Orozco_2019_KernelDBS], which approach VTA in different ways.
+The Gaussian white-noise model is adopted because it furnishes the maximally uninformative representation of uncertainty under a finite-variance constraint [Jaynes2003, CoverThomas2006].
+Given that the present optimization framework does not specify a detailed probabilistic law for perturbations, this choice avoids introducing unsupported structure and provides a neutral, information-theoretic baseline.
+More sophisticated uncertainty quantification techniques, such as Markov Chain Monte Carlo (MCMC)-based Bayesian sampling [kaipio2006statistical] or deep-learning approaches to predictive uncertainty [schmidt2016uncertainty, athawale2019statistical], could be employed when richer data are available.
+Compared to the reference methods, the L1L1 method is also slightly more demanding in terms of computing time.
 However, the increase is not significant and does not limit its practical use.
-Other mathematical optimization techniques, such as Tikhonov (TLS) [38, 37], may offer promising alternatives alongside the L1L1 method for further refining the a priori uncertainty modeling adopted in this study.
+Other mathematical optimization techniques, such as TLS, may offer promising alternatives alongside the L1L1 method for further refining the a priori uncertainty modeling adopted in this study.
 A logical direction for future research is the development of a posteriori models, where the current approach could be validated in more advanced experimental or clinical settings.
@@
 Translating these computational findings into clinical practice will require rigorous validation through empirical testing and clinical trials.
-An ongoing work is the inclusion of more patient-specific anatomical and physiological data, e.g., a thalamic segmentation, to ensure more personalized and accurate stimulation strategies.
-Therefore, we explore the use of FE-based discretizations incorporating thalamic subdivisions via boundary-fitted multicompartment FE meshing approach and a complete electrode model, as utilized in this study.
+This manuscript does not address clinical implications; however, for the configuration studied, certain anatomical structures, such as the lamina medullaris near the modeled lead trajectory, may require special consideration in clinical practice to avoid stimulation-induced side effects.
+Translating these methodological results to clinical applications will necessitate examining the relationship between the L1L1 method and uncertainties in individual factors such as lead placement, tissue conductivity and anisotropy, and patient-specific anatomical variations including thalamic segmentation.
+To support this, we employ finite element (FE) discretizations incorporating thalamic subdivisions through a boundary-fitted multi-compartment FE mesh together with a Complete Electrode Model (CEM) [pursiainen2017forward], as in the present study.
 5 Conclusions
 This study represents an advancement in optimizing the volumetric current fields for DBS by leveraging the metaheuristic L1L1 method.
-Previous studies on contact selection have used convex optimization and patient-specific models to determine optimal stimulation settings.
-These methods perform well when all input parameters are fixed.
-However, they do not account for uncertainties in factors such as patient anatomy, lead placement, and tissue properties, which often vary in clinical practice.
-In principle, our L1L1-based method can address these uncertainties.
-Its main advantage is the ability to target the desired area while avoiding overfitting to noisy data.
-Our findings demonstrate the method’s ability to refine electrode contact point configurations under varying levels of a priori lead field uncertainty.
-By achieving relatively high field ratios with a carefully chosen $\varepsilon$ value, the L1L1 method shows promise in enhancing the precision and effectiveness of DBS stimulation within uncertainty constraints.
+Previous contact selection studies, such as [Anderson_2018] and later [Janson_2020_Activation], have used convex optimization to determine optimal stimulation settings and this represents at addressing varying levels of a priori lead field uncertainty, which can be attributed as patient anatomy, misleading lead placement, and tissue properties, which often vary in clinical practice.
+Our findings demonstrate that the method can achieve relatively high field ratios with a carefully chosen activation threshold $\varepsilon$ value.
 This, in turn, has the potential to improve individualized stimulation outcomes.
 An ongoing work is to integrate patient-specific data, such as thalamic segmentation, to improve stimulation accuracy.
-This involves using boundary-fitted multicompartment FE meshes and a complete electrode model, as in this study.
+This involves using boundary-fitted multi-compartment FE meshes [galazprieto_2023_mesh] and a Complete Electrode Model [pursiainen2017forward], as in this study.
 6 Conflict of Interest
@@
 7 Author Contributions
-FGP: Conceptualization, Formal analysis, Investigation, Methodology, Software, Visualization, Writing—original draft.
+FGP: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Visualization, Writing—original draft.
 AL: Conceptualization, Data curation, Investigation, Methodology, Project administration, Writing—original draft.
@@
 8 Funding
-FGP, MS, and SP were supported by the Research Council of Finland through the Center of Excellence in Inverse Modelling and Imaging 2018–2025 (353089); the researcher exchange (DAAD) project ”Non-invasively reconstructing and inhibiting activity in focal epilepsy” (354976, 367453); the ERA PerEpi project (344712), and Flagship of Advanced Mathematics for Sensing, Imaging and Modelling (FAME) (359185, 359198)
+FGP, MS, and SP were supported by the Research Council of Finland (RCF) through the Center of Excellence in Inverse Modelling and Imaging 2018–2025 (RCF 353089), Exploratory Study for Radar Tomography of Dimorphos (RCF 359198), and the Flagship of Advanced Mathematics for Sensing, Imaging and Modelling (FAME) (RCF 359185); DAAD (German Academic Exchange Service) project ”Non-invasively reconstructing and inhibiting activity in focal epilepsy” (354976, 367453); and the ERA Personalised diagnosis and treatment for refractory focal paediatric and adult epilepsy (PerEpi) project (RCF 344712).
+MS holds Academy Research Fellowship 2025–2029 (RCF 371055).
+9 Ethical Statement
+This work is based exclusively on mathematical modeling and numerical simulations, using an openly available MRI dataset [piastra_maria_carla_2020_3888381] as input.
+No new data was collected, and no sensitive or personally identifiable information was processed at any stage of the study.
 10 Acknowledgments
@@
 Appendix A Mathematical Optimization Scheme
-We consider a weighted optimization scheme governed by the linear system
-${\bf L}{\bf y}={\bf x},$ (1)
-where $\mathbf{y}\in\mathbb{R}^{K}$ is the current pattern to be optimized, $\mathbf{x}\in\mathbb{R}^{N}$ is the discretized volume current density vector, and $\mathbf{L}\in\mathbb{R}^{N\times K}$ is the lead field matrix defining a linear forward mapping from current patterns to the resulting current density field.
-This system is derived by component-wise splitting of the full lead field matrix $\hat{\mathbf{L}}\in\mathbb{R}^{N\times K}$ and the full current field $\hat{\mathbf{x}}\in\mathbb{R}^{N}$, such that
-$\hat{\mathbf{L}}=\begin{pmatrix}\hat{\mathbf{L}}_{1}\\ \hat{\mathbf{L}}_{2}\end{pmatrix},\quad\hat{\mathbf{x}}=\begin{pmatrix}\hat{\mathbf{x}}_{1}\\ \mathbf{0}\end{pmatrix},$ (2)
-where $\hat{\mathbf{x}}_{1}\in\mathbb{R}^{N_{1}}$ corresponds to the target region of interest and the zero component enforces field suppression elsewhere. The reduced system (1) is then defined by
-$\mathbf{L}=\begin{pmatrix}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{pmatrix}=\begin{pmatrix}\mathbf{P}\hat{\mathbf{L}}_{1}\\ \hat{\mathbf{L}}_{2}\end{pmatrix},\quad\mathbf{x}=\begin{pmatrix}\mathbf{x}_{1}\\ \mathbf{0}\end{pmatrix}=\begin{pmatrix}\mathbf{P}\hat{\mathbf{x}}_{1}\\ \mathbf{0}\end{pmatrix}.$ (3)
-Here, $\mathbf{P}\in\mathbb{R}^{N_{1}\times N_{1}}$ is a projection matrix that projects vectors onto the direction of $\hat{\mathbf{x}}_{1}$, which is assumed to have fixed direction.
-This projection ensures that alignment between the induced and target fields is measured only along the direction of interest.
-In this formulation, the upper part $\hat{\mathbf{L}}_{1}\mathbf{y}=\hat{\mathbf{x}}_{1}$ defines the focused current field in the target region, while the lower part $\hat{\mathbf{L}}_{2}\mathbf{y}=\mathbf{0}$ defines the nuisance field to be suppressed outside the region of interest.
-A.1 L1-norm regularized L1-norm fitting
+A.1 Formulation of the Lead Field Matrix
+The linear forward mapping between electrode currents and the resulting volumetric current density distribution is given by the real-valued $(N+M)\times K$ lead field matrix $\mathbf{L}$:
+$\mathbf{L}\,\mathbf{y}=\mathbf{x},$ (1)
+where $\mathbf{y}\in\mathbb{R}^{K}$ denotes the applied current pattern across $K$ electrode contacts, and $\mathbf{x}\in\mathbb{R}^{N+M}$ is the discretized volumetric current density with each entry corresponding to a local source element parameterized by three Cartesian dipolar degrees of freedom (DOFs), representing current density along the $x$-, $y$-, and $z$-axes.
+For our optimization framework, $\mathbf{L}$ is partitioned into target and nuisance components:
+$\mathbf{L}=\begin{pmatrix}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{pmatrix},\qquad\mathbf{x}=\begin{pmatrix}\mathbf{x}_{1}\\ \mathbf{0}\end{pmatrix}$, (2)
+where $\mathbf{L}_{1}\in\mathbb{R}^{N\times K}$ models the target region, $\mathbf{L}_{2}\in\mathbb{R}^{M\times K}$ models the nuisance region, and $\mathbf{x}_{1}\in\mathbb{R}^{N}$ defines the desired field in the target region.
+To ensure that only the direction of $\mathbf{x}_{1}$ is prescribed, we introduce the projection
+$\mathbf{P}=\frac{\mathbf{x}_{1}\mathbf{x}_{1}^{\mathsf{T}}}{\|\mathbf{x}_{1}\|_{2}^{2}}\in\mathbb{R}^{N\times N},$ (3)
+which enforces alignment of the induced field with $\mathbf{x}_{1}$. The resulting reduced formulation is:
+$\begin{pmatrix}\mathbf{P}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{pmatrix}\mathbf{y}=\begin{pmatrix}\mathbf{P}\mathbf{x}_{1}\\ \mathbf{0}\end{pmatrix}$. (4)
+The first block enforces field orientation within the target region, while the second enforces suppression outside it.
+A.2 L1-norm Regularized L1-norm Fitting (L1L1)
 The general form of the L1-norm regularized L1-norm fitting (L1L1) method described by Galaz Prieto et.
-al [32, 57] is an optimization problem designed to minimize a linear objective function of continuous real variables subject to linear constraints.
+al [galazprieto_2022_L1vsL2] is a linear programming optimization problem designed to minimize a linear objective function of continuous real variables subject to linear constraints.
 The optimization problem needs to find the best matching between $\bf y$}, and the focused field via ${\bf L}{\bf y}={\bf x}$.
 The task is
-[…]
+$\underset{y}{\textrm{minimize}}$ $\quad\left\|\begin{pmatrix}{\bf L}_{1}{\bf y}-{\bf x}_{1}\\ \Psi_{\varepsilon}[\nu^{-1}{\bf L}_{2}{\bf y}]\end{pmatrix}\right\|_{1}+{\alpha}\zeta\|{\bf y}\|_{1},$ (5)
 subject to $\quad{\bf y}\preceq\gamma{\bf 1},$ (6)
@@
 $\quad\sum_{\ell=1}^{K}y_{\ell}=0,\,\quad\ell=1,\ldots,K.$ (8)
-The linear constraint 5 implies that the individual current injection for each $\ell$-th electrode is limited to a $\Gamma$ value; constraint 6 implies the total current injection flowing through the system is within a safety $\mu$-value limit; and constraint 7 indicates the total sum of electrical activity from every active electrode in the system must be equal to zero.
+The linear constraint 6 implies that the individual current injection for each $\ell$-th electrode is limited to a $\gamma$ value; constraint 7 implies the total current injection flowing through the system is within a safety $\mu$-value limit; and constraint 8 indicates the total sum of electrical activity from every active electrode in the system must be equal to zero.
 The regularization parameter $\bf\alpha$} sets the L1-regularization level regarding a scaling value $\zeta=\|{\bf L}\|_{1}$.
@@
 The constraint support index $\{\,m\,:\,|({\bf L}_{2}{\bf y})_{m}|\geq{\varepsilon\nu}\,\}$, i.e., sets the contributing to the value of the objective function.
-A.2 Tikhonov-Regularized Least Squares
-In TLS estimation [37, 38] is formulated as the following optimization problem:
+To evaluate stimulation performance and to quantify the trade-off between target and undesired current spread, we define three decision variables: the focused current density $\Gamma$, the nuisance current density $\Xi$, and the resulting field ratio $\Theta$.
+The focused current density $\Gamma$ (A/m2) quantifies the degree of field alignment between the induced current distribution and the desired stimulation target. It is defined as
+$\Gamma=\frac{{\bf x}_{1}^{T}{\bf L}_{1}{\bf y}}{\|{\bf x}_{1}\|_{2}},$ (10)
+where ${\bf x}_{1}\in\mathbb{R}^{N}$ represents the target volumetric current density template specifying the spatial and directional characteristics of the desired field within the stimulation region.
+The inner product ${\bf x}_{1}^{T}{\bf L}_{1}{\bf y}$ measures how strongly the induced field ${\bf L}_{1}{\bf y}$ aligns with the target pattern, while normalization by $\|{\bf x}_{1}\|_{2}$ scales the result to yield a field intensity expressed in A/m2.
+The nuisance current density $\Xi$ (A/m2) characterizes the average magnitude of the induced field in off-target regions and is defined by
+$\Xi=\frac{\|{\bf L}_{2}{\bf y}\|_{2}}{\sqrt{M}}.$ (11)
+Dividing the Euclidean norm $\|{\bf L}_{2}{\bf y}\|_{2}$ by $\sqrt{M}$ yields the root-mean-square (RMS) field amplitude per node,
+$\frac{\|{\bf L}_{2}{\bf y}\|_{2}}{\sqrt{M}}=\sqrt{\frac{1}{M}\sum_{m=1}^{M}\big(({\bf L}_{2}{\bf y})_{m}\big)^{2}}.$
+The division by $\sqrt{M}$ converts the Euclidean norm into the root-mean-square (RMS) field amplitude per node, thereby making $\Xi$ invariant to the total number of discretization points in the nuisance domain.
+In cases where only a subset of nodes exhibits appreciable current density, the normalization may alternatively be restricted to the active constraint support, i.e., replacing $M$ with the number of nodes satisfying $|({\bf L}_{2}{\bf y})_{m}|\geq\varepsilon\nu$.
+The field ratio $\Theta$ is introduced to characterize the overall selectivity of stimulation:
+$\Theta=\frac{\Gamma}{\Xi}.$ (12)
+A high $\Theta$ value reflects both strong current focusing within the target and minimal activation of non-target regions.
+A.3 Tikhonov-Regularized Least Squares
+TLS estimation [dmochowski2011optimized] is formulated as the following optimization problem:
 $\min_{{\bf y}}\left\{\|{\bf L}_{1}{\bf y}-{\bf x}_{1}\|^{2}_{2}+\gamma^{2}\beta^{2}\|{\bf L}_{2}{\bf y}\|^{2}_{2}+\gamma^{2}\varsigma^{2}\|{\bf y}\|^{2}_{2}\right\},$ (13)
@@
 Both the regularization parameter and the nuisance field weight influence the above equations.
-However, different methods handle the latter distinctly: in L1L1, it appears as an explicit threshold, whereas in the TLS framework, it is incorporated as a weighted term, leading to a distinct dependence of the solution.
-For a more detailed description of the methods above, forward model and comparative results using other optimization approaches such as the L1-norm regularized L2-norm fitting (L1L2) method, a variant of the L1L1 method, please consult the [32] paper.
-For technicalities and explanation of the mathematical optimization packages, algorithms and search methodologies applied, see the follow-up [57] paper.